GNU Automake

For version 1.7.1, 25 September 2002

David MacKenzie and Tom Tromey

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995, 1996, 2000, 2001, 2002 Free Software Foundation, Inc.

This is the first edition of the GNU Automake documentation, and is consistent with GNU Automake 1.7.1.

Published by the Free Software Foundation 59 Temple Place - Suite 330, Boston, MA 02111-1307 USA

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

1 Introduction

Automake is a tool for automatically generating Makefile.ins from files called Makefile.am. Each Makefile.am is basically a series of make variable definitions^{1}, with rules being thrown in occasionally. The generated Makefile.ins are compliant with the GNU Makefile standards.

The GNU Makefile Standards Document (see Section “Makefile Conventions” in The GNU Coding Standards) is long, complicated, and subject to change. The goal of Automake is to remove the burden of Makefile maintenance from the back of the individual GNU maintainer (and put it on the back of the Automake maintainer).

The typical Automake input file is simply a series of variable definitions. Each such file is processed to create a Makefile.in. There should generally be one Makefile.am per directory of a project.

Automake does constrain a project in certain ways; for instance it assumes that the project uses Autoconf (see Section “Introduction” in The Autoconf Manual), and enforces certain restrictions on the configure.in contents^{2}.

Automake requires perl in order to generate the Makefile.ins. However, the distributions created by Automake are fully GNU standards-compliant, and do not require perl in order to be built.

Mail suggestions and bug reports for Automake to bug-automake@gnu.org.

2 General ideas

The following sections cover a few basic ideas that will help you understand how Automake works.

2.1 General Operation

Automake works by reading a Makefile.am and generating a Makefile.in. Certain variables and targets defined in the Makefile.am instruct Automake to generate more specialized code; for instance, a ‘bin_PROGRAMS’ variable definition will cause targets for compiling and linking programs to be generated.

The variable definitions and targets in the Makefile.am are copied verbatim into the generated file. This allows you to add arbitrary code into the generated Makefile.in. For instance the Automake distribution includes a non-standard cvs-dist target, which the Automake maintainer uses to make distributions from his source control system.

Note that most GNU make extensions are not recognized by Automake. Using such extensions in a Makefile.am will lead to errors or confusing behavior.

A special exception is that the GNU make append operator, ‘+=’, is supported. This operator appends its right hand argument to the variable specified on the left. Automake will translate the operator into an ordinary ‘=’ operator; ‘+=’ will thus work with any make program.

1 These variables are also called make macros in Make terminology, however in this manual we reserve the

term macro for Autoconf’s macros.

2 Autoconf 2.50 promotes configure.ac over configure.in. The rest of this documentation will refer to

configure.in as this use is not yet spread, but Automake supports configure.ac too.

Automake tries to keep comments grouped with any adjoining targets or variable definitions.

A target defined in Makefile.am generally overrides any such target of a similar name that would be automatically generated by automake. Although this is a supported feature, it is generally best to avoid making use of it, as sometimes the generated rules are very particular.

Similarly, a variable defined in Makefile.am or AC_SUBST’ed from configure.in will override any definition of the variable that automake would ordinarily create. This feature is more often useful than the ability to override a target definition. Be warned that many of the variables generated by automake are considered to be for internal use only, and their names might change in future releases.

When examining a variable definition, Automake will recursively examine variables referenced in the definition. For example, if Automake is looking at the content of foo_SOURCES in this snippet

     xs = a.c b.c
     foo_SOURCES = c.c $(xs)

it would use the files a.c, b.c, and c.c as the contents of foo_SOURCES.

Automake also allows a form of comment which is not copied into the output; all lines beginning with ‘##’ (leading spaces allowed) are completely ignored by Automake.

It is customary to make the first line of Makefile.am read:

     ## Process this file with automake to produce Makefile.in

2.2 Strictness

While Automake is intended to be used by maintainers of GNU packages, it does make some effort to accommodate those who wish to use it, but do not want to use all the GNU conventions.

To this end, Automake supports three levels of strictness—the strictness indicating how stringently Automake should check standards conformance.

The valid strictness levels are:

‘foreign’ Automake will check for only those things which are absolutely required for proper operations. For instance, whereas GNU standards dictate the existence of a NEWS file, it will not be required in this mode. The name comes from the fact that Automake is intended to be used for GNU programs; these relaxed rules are not the standard mode of operation.

‘gnu’ Automake will check—as much as possible—for compliance to the GNU standards for packages. This is the default.

‘gnits’ Automake will check for compliance to the as-yet-unwritten Gnits standards. These are based on the GNU standards, but are even more detailed. Unless you are a Gnits standards contributor, it is recommended that you avoid this option until such time as the Gnits standard is actually published (which may never happen).

For more information on the precise implications of the strictness level, see Chapter 21 [Gnits], page 57.

Automake also has a special “cygnus” mode which is similar to strictness but handled differently. This mode is useful for packages which are put into a “Cygnus” style tree (e.g., the GCC tree). For more information on this mode, see Chapter 22 [Cygnus], page 58.

2.3 The Uniform Naming Scheme

Automake variables generally follow a uniform naming scheme that makes it easy to decide how programs (and other derived objects) are built, and how they are installed. This scheme also supports configure time determination of what should be built.

At make time, certain variables are used to determine which objects are to be built. The variable names are made of several pieces which are concatenated together.

The piece which tells automake what is being built is commonly called the primary. For instance, the primary PROGRAMS holds a list of programs which are to be compiled and linked.

A different set of names is used to decide where the built objects should be installed. These names are prefixes to the primary which indicate which standard directory should be used as the installation directory. The standard directory names are given in the GNU standards (see Section “Directory Variables” in The GNU Coding Standards). Automake extends this list with pkglibdir, pkgincludedir, and pkgdatadir; these are the same as the non-‘pkg’ versions, but with ‘@PACKAGE@’ appended. For instance, pkglibdir is defined

as $(libdir)/@PACKAGE@.

For each primary, there is one additional variable named by prepending ‘EXTRA_’ to the primary name. This variable is used to list objects which may or may not be built, depending on what configure decides. This variable is required because Automake must statically know the entire list of objects that may be built in order to generate a Makefile.in that will work in all cases.

For instance, cpio decides at configure time which programs are built. Some of the programs are installed in bindir, and some are installed in sbindir:

     EXTRA_PROGRAMS = mt rmt
     bin_PROGRAMS = cpio pax
     sbin_PROGRAMS = @MORE_PROGRAMS@

Defining a primary without a prefix as a variable, e.g., PROGRAMS, is an error.

Note that the common ‘dir’ suffix is left off when constructing the variable names; thus

one writes ‘bin_PROGRAMS’ and not ‘bindir_PROGRAMS’.

Not every sort of object can be installed in every directory. Automake will flag those attempts it finds in error. Automake will also diagnose obvious misspellings in directory names.

Sometimes the standard directories—even as augmented by Automake— are not enough. In particular it is sometimes useful, for clarity, to install objects in a subdirectory of some predefined directory. To this end, Automake allows you to extend the list of possible installation directories. A given prefix (e.g. ‘zar’) is valid if a variable of the same name with ‘dir’ appended is defined (e.g. zardir).

For instance, until HTML support is part of Automake, you could use this to install raw HTML documentation:

     htmldir = $(prefix)/html
     html_DATA = automake.html

The special prefix ‘noinst’ indicates that the objects in question should be built but not installed at all. This is usually used for objects required to build the rest of your package, for instance static libraries (see Section 9.2 [A Library], page 26), or helper scripts.

The special prefix ‘check’ indicates that the objects in question should not be built until the make check command is run. Those objects are not installed either.

The current primary names are ‘PROGRAMS’, ‘LIBRARIES’, ‘LISP’, ‘PYTHON’, ‘JAVA’,

‘SCRIPTS’, ‘DATA’, ‘HEADERS’, ‘MANS’, and ‘TEXINFOS’.

Some primaries also allow additional prefixes which control other aspects of automake’s behavior. The currently defined prefixes are ‘dist_’, ‘nodist_’, and ‘nobase_’. These prefixes are explained later (see Section 9.4 [Program and Library Variables], page 27).

2.4 How derived variables are named

Sometimes a Makefile variable name is derived from some text the maintainer supplies. For instance, a program name listed in ‘_PROGRAMS’ is rewritten into the name of a ‘_SOURCES’ variable. In cases like this, Automake canonicalizes the text, so that program names and the like do not have to follow Makefile variable naming rules. All characters in the name except for letters, numbers, the strudel (@), and the underscore are turned into underscores when making variable references.

For example, if your program is named sniff-glue, the derived variable name would be sniff_glue_SOURCES, not sniff-glue_SOURCES. Similarly the sources for a library named

libmumble++.a should be listed in the libmumble___a_SOURCES variable.

The strudel is an addition, to make the use of Autoconf substitutions in variable names less obfuscating.

2.5 Variables reserved for the user

Some Makefile variables are reserved by the GNU Coding Standards for the use of the “user” – the person building the package. For instance, CFLAGS is one such variable.

Sometimes package developers are tempted to set user variables such as CFLAGS because it appears to make their job easier – they don’t have to introduce a second variable into every target.

However, the package itself should never set a user variable, particularly not to include switches which are required for proper compilation of the package. Since these variables are documented as being for the package builder, that person rightfully expects to be able to override any of these variables at build time.

To get around this problem, automake introduces an automake-specific shadow variable for each user flag variable. (Shadow variables are not introduced for variables like CC, where they would make no sense.) The shadow variable is named by prepending ‘AM_’ to the user variable’s name. For instance, the shadow variable for YFLAGS is AM_YFLAGS.

2.6 Programs automake might require

Automake sometimes requires helper programs so that the generated Makefile can do its work properly. There are a fairly large number of them, and we list them here.

ansi2knr.c
ansi2knr.1

These two files are used by the automatic de-ANSI-fication support (see Section 9.13 [ANSI], page 38).

compile This is a wrapper for compilers which don’t accept both ‘-c’ and ‘-o’ at the same time. It is only used when absolutely required. Such compilers are rare.

config.guess
config.sub

These programs compute the canonical triplets for the given build, host, or target architecture. These programs are updated regulary to support new architectures and fix probes broken by changes in new kernel versions. You are encouraged to fetch the latest versions of these files from ftp://ftp.gnu.org/ gnu/config/ before making a release.

depcomp This program understands how to run a compiler so that it will generate not only the desired output but also dependency information which is then used by the automatic dependency tracking feature.

elisp-comp

This program is used to byte-compile Emacs Lisp code.

install-sh

This is a replacement for the install program which works on platforms where install is unavailable or unusable.

mdate-sh This script is used to generate a version.texi file. It examines a file and prints some date information about it.

missing This wraps a number of programs which are typically only required by maintainers. If the program in question doesn’t exist, missing prints an informative warning and attempts to fix things so that the build can continue.

mkinstalldirs

This works around the fact that mkdir -p is not portable.

py-compile

This is used to byte-compile Python scripts.

texinfo.tex

Not a program, this file is required for make dvi, make ps and make pdf to work when Texinfo sources are in the package.

ylwrap This program wraps lex and yacc and ensures that, for instance, multiple yacc instances can be invoked in a single directory in parallel.

3 Some example packages

3.1 A simple example, start to finish

Let’s suppose you just finished writing zardoz, a program to make your head float from vortex to vortex. You’ve been using Autoconf to provide a portability framework, but your Makefile.ins have been ad-hoc. You want to make them bulletproof, so you turn to Automake.

The first step is to update your configure.in to include the commands that automake needs. The way to do this is to add an AM_INIT_AUTOMAKE call just after AC_INIT:

     AC_INIT(zardoz, 1.0)
     AM_INIT_AUTOMAKE
     ...

Since your program doesn’t have any complicating factors (e.g., it doesn’t use gettext, it doesn’t want to build a shared library), you’re done with this part. That was easy!

Now you must regenerate configure. But to do that, you’ll need to tell autoconf how to find the new macro you’ve used. The easiest way to do this is to use the aclocal program to generate your aclocal.m4 for you. But wait… maybe you already have an aclocal.m4, because you had to write some hairy macros for your program. The aclocal program lets you put your own macros into acinclude.m4, so simply rename and then run:

     mv aclocal.m4 acinclude.m4
     aclocal
     autoconf

Now it is time to write your Makefile.am for zardoz. Since zardoz is a user program, you want to install it where the rest of the user programs go: bindir. Additionally, zardoz has some Texinfo documentation. Your configure.in script uses AC_REPLACE_FUNCS, so you need to link against ‘@LIBOBJS@’. So here’s what you’d write:

     bin_PROGRAMS = zardoz
     zardoz_SOURCES = main.c head.c float.c vortex9.c gun.c
     zardoz_LDADD = @LIBOBJS@
     info_TEXINFOS = zardoz.texi

Now you can run automake --add-missing to generate your Makefile.in and grab any auxiliary files you might need, and you’re done!

3.2 A classic program

GNU hello (ftp://prep.ai.mit.edu/pub/gnu/hello-1.3.tar.gz) is renowned for its

classic simplicity and versatility. This section shows how Automake could be used with the GNU Hello package. The examples below are from the latest beta version of GNU Hello, but with all of the maintainer-only code stripped out, as well as all copyright comments.

Of course, GNU Hello is somewhat more featureful than your traditional two-liner. GNU Hello is internationalized, does option processing, and has a manual and a test suite.

Here is the configure.in from GNU Hello:

     dnl Process this file with autoconf to produce a configure script.

AC_INIT(src/hello.c)
AM_INIT_AUTOMAKE(hello, 1.3.11)
AM_CONFIG_HEADER(config.h)
dnl Set of available languages.
ALL_LINGUAS="de fr es ko nl no pl pt sl sv"
dnl Checks for programs.
AC_PROG_CC
AC_ISC_POSIX
dnl Checks for libraries.
dnl Checks for header files.
AC_STDC_HEADERS
AC_HAVE_HEADERS(string.h fcntl.h sys/file.h sys/param.h)
dnl Checks for library functions.
AC_FUNC_ALLOCA
dnl Check for st_blksize in struct stat
AC_ST_BLKSIZE
dnl internationalization macros
AM_GNU_GETTEXT
AC_OUTPUT([Makefile doc/Makefile intl/Makefile po/Makefile.in \
           src/Makefile tests/Makefile tests/hello],
   [chmod +x tests/hello])

The ‘AM_’ macros are provided by Automake (or the Gettext library); the rest are standard Autoconf macros.

The top-level Makefile.am:

EXTRA_DIST = BUGS ChangeLog.O
SUBDIRS = doc intl po src tests

As you can see, all the work here is really done in subdirectories. The po and intl directories are automatically generated using gettextize; they will

not be discussed here.

In doc/Makefile.am we see:
info_TEXINFOS = hello.texi
hello_TEXINFOS = gpl.texi

This is sufficient to build, install, and distribute the GNU Hello manual.

Here is tests/Makefile.am:
TESTS = hello
EXTRA_DIST = hello.in testdata

The script hello is generated by configure, and is the only test case. make check will run this test.

Last we have src/Makefile.am, where all the real work is done:

     bin_PROGRAMS = hello
     hello_SOURCES = hello.c version.c getopt.c getopt1.c getopt.h system.h
     hello_LDADD = @INTLLIBS@ @ALLOCA@
     localedir = $(datadir)/locale
     INCLUDES = -I../intl -DLOCALEDIR=\"$(localedir)\"

3.3 Building true and false

Here is another, trickier example. It shows how to generate two programs (true and false) from the same source file (true.c). The difficult part is that each compilation of true.c requires different cpp flags.

     bin_PROGRAMS = true false
     false_SOURCES =
     false_LDADD = false.o
     true.o: true.c
             $(COMPILE) -DEXIT_CODE=0 -c true.c
     false.o: true.c
             $(COMPILE) -DEXIT_CODE=1 -o false.o -c true.c

Note that there is no true_SOURCES definition. Automake will implicitly assume that there is a source file named true.c, and define rules to compile true.o and link true. The true.o: true.c rule supplied by the above Makefile.am, will override the Automake generated rule to build true.o.

false_SOURCES is defined to be empty—that way no implicit value is substituted. Because we have not listed the source of false, we have to tell Automake how to link the program. This is the purpose of the false_LDADD line. A false_DEPENDENCIES variable, holding the dependencies of the false target will be automatically generated by Automake from the contant of false_LDADD.

The above rules won’t work if your compiler doesn’t accept both ‘-c’ and ‘-o’. The simplest fix for this is to introduce a bogus dependency (to avoid problems with a parallel

make):
     true.o: true.c false.o
             $(COMPILE) -DEXIT_CODE=0 -c true.c
     false.o: true.c
             $(COMPILE) -DEXIT_CODE=1 -c true.c && mv true.o false.o

Also, these explicit rules do not work if the de-ANSI-fication feature is used (see Section 9.13 [ANSI], page 38). Supporting de-ANSI-fication requires a little more work:

     true._o: true._c false.o
             $(COMPILE) -DEXIT_CODE=0 -c true.c
     false._o: true._c
             $(COMPILE) -DEXIT_CODE=1 -c true.c && mv true._o false.o

As it turns out, there is also a much easier way to do this same task. Some of the above techniques are useful enough that we’ve kept the example in the manual. However if you were to build true and false in real life, you would probably use per-program compilation flags, like so:

     bin_PROGRAMS = false true
     false_SOURCES = true.c
     false_CPPFLAGS = -DEXIT_CODE=1
     true_SOURCES = true.c
     true_CPPFLAGS = -DEXIT_CODE=0

In this case Automake will cause true.c to be compiled twice, with different flags. De-ANSI-fication will work automatically. In this instance, the names of the object files would be chosen by automake; they would be false-true.o and true-true.o. (The name of the object files rarely matters.)

4 Creating a Makefile.in

To create all the Makefile.ins for a package, run the automake program in the top level directory, with no arguments. automake will automatically find each appropriate Makefile.am (by scanning configure.in; see Chapter 5 [configure], page 12) and generate the corresponding Makefile.in. Note that automake has a rather simplistic view of what constitutes a package; it assumes that a package has only one configure.in, at the top. If your package has multiple configure.ins, then you must run automake in each directory holding a configure.in. (Alteratively, you may rely on Autoconf’s autoreconf, which is able to recurse your package tree and run automake where appropriate.)

You can optionally give automake an argument; .am is appended to the argument and the result is used as the name of the input file. This feature is generally only used to automatically rebuild an out-of-date Makefile.in. Note that automake must always be run from the topmost directory of a project, even if being used to regenerate the Makefile.in in some subdirectory. This is necessary because automake must scan configure.in, and because automake uses the knowledge that a Makefile.in is in a subdirectory to change its behavior in some cases.

Automake will run autoconf to scan configure.in and its dependencies (aclocal.m4), therefore autoconf must be in your PATH. If there is an AUTOCONF variable in your environment it will be used instead of autoconf, this allows you to select a particular version of Autoconf. By the way, don’t misunderstand this paragraph: Automake runs autoconf to scan your configure.in, this won’t build configure and you still have to run autoconf yourself for this purpose.

automake accepts the following options:

‘-a’
‘--add-missing’

Automake requires certain common files to exist in certain situations; for in-

          stance config.guess is required if configure.in runs AC_CANONICAL_HOST.

Automake is distributed with several of these files (see Section 2.6 [Auxiliary Programs], page 5); this option will cause the missing ones to be automatically added to the package, whenever possible. In general if Automake tells you a file is missing, try using this option. By default Automake tries to make a symbolic link pointing to its own copy of the missing file; this can be changed

with --copy.
‘--libdir=dir’

Look for Automake data files in directory dir instead of in the installation directory. This is typically used for debugging.

‘-c’

‘--copy’ When used with --add-missing, causes installed files to be copied. The default is to make a symbolic link.

‘--cygnus’

Causes the generated Makefile.ins to follow Cygnus rules, instead of GNU or Gnits rules. For more information, see Chapter 22 [Cygnus], page 58.

‘-f’
‘--force-missing’

When used with --add-missing, causes standard files to be reinstalled even if they already exist in the source tree. This involves removing the file from the source tree before creating the new symlink (or, with --copy, copying the new file).

‘--foreign’

Set the global strictness to ‘foreign’. For more information, see Section 2.2 [Strictness], page 2.

‘--gnits’ Set the global strictness to ‘gnits’. For more information, see Chapter 21 [Gnits], page 57.

‘--gnu’ Set the global strictness to ‘gnu’. For more information, see Chapter 21 [Gnits], page 57. This is the default strictness.

‘--help’ Print a summary of the command line options and exit.

‘-i’
‘--ignore-deps’

This disables the dependency tracking feature in generated Makefiles; see Section 9.14 [Dependencies], page 39.

‘--include-deps’

This enables the dependency tracking feature. This feature is enabled by default. This option is provided for historical reasons only and probably should not be used.

‘--no-force’

Ordinarily automake creates all Makefile.ins mentioned in configure.in. This option causes it to only update those Makefile.ins which are out of date with respect to one of their dependents.

‘-o dir’
‘--output-dir=dir’

Put the generated Makefile.in in the directory dir. Ordinarily each Makefile.in is created in the directory of the corresponding Makefile.am. This option is deprecated and will be removed in a future release.

‘-v’
‘--verbose’

Cause Automake to print information about which files are being read or created.

‘--version’

Print the version number of Automake and exit.

‘-W CATEGORY’
‘--warnings=category’

Output warnings falling in category. category can be one of:

‘gnu’ warnings related to the GNU Coding Standards (see The GNU Coding Standards).

‘obsolete’

obsolete features or constructions

‘portability’

portability issues (e.g., use of Make features which are known not portable)

‘syntax’ weird syntax, unused variables, typos

‘unsupported’

unsupported or incomplete features

‘all’ all the warnings

‘none’ turn off all the warnings

‘error’ treat warnings as errors

A category can be turned off by prefixing its name with ‘no-’. For instance ‘-Wno-syntax’ will hide the warnings about unused variables.

The categories output by default are ‘syntax’ and ‘unsupported’. Additionally, ‘gnu’ is enabled in ‘--gnu’ and ‘--gnits’ strictness.

‘portability’ warnings are currently disabled by default, but they will be enabled in ‘--gnu’ and ‘--gnits’ strictness in a future release.

The environment variable ‘WARNINGS’ can contain a comma separated list of categories to enable. It will be taken into account before the command-line switches, this way ‘-Wnone’ will also ignore any warning category enabled by ‘WARNINGS’. This variable is also used by other tools like autoconf; unknown categories are ignored for this reason.

5 Scanning configure.in

Automake scans the package’s configure.in to determine certain information about the package. Some autoconf macros are required and some variables must be defined in configure.in. Automake will also use information from configure.in to further tailor its output.

Automake also supplies some Autoconf macros to make the maintenance easier. These macros can automatically be put into your aclocal.m4 using the aclocal program.

5.1 Configuration requirements

The one real requirement of Automake is that your configure.in call AM_INIT_AUTOMAKE. This macro does several things which are required for proper Automake operation (see Section 5.6 [Macros], page 17).

Here are the other macros which Automake requires but which are not run by AM_INIT_

AUTOMAKE:
AC_CONFIG_FILES
AC_OUTPUT

Automake uses these to determine which files to create (see Section “Creating Output Files” in The Autoconf Manual). A listed file is considered to be an Automake generated Makefile if there exists a file with the same name and the

          .am extension appended. Typically, AC_CONFIG_FILES([foo/Makefile]) will

cause Automake to generate foo/Makefile.in if foo/Makefile.am exists.

Other listed files are treated differently. Currently the only difference is that an Automake Makefile is removed by make distclean, while other files are removed by make clean.

5.2 Other things Automake recognizes

Every time Automake is run it calls Autoconf to trace configure.in. This way it can recognize the use of certain macros and tailor the generated Makefile.in appropriately. Currently recognized macros and their effects are:

AC_CONFIG_HEADERS

Automake will generate rules to rebuild these headers. Older versions of Automake required the use of AM_CONFIG_HEADER (see Section 5.6 [Macros], page 17); this is no longer the case today.

AC_CONFIG_AUX_DIR

Automake will look for various helper scripts, such as mkinstalldirs, in the directory named in this macro invocation. If not seen, the scripts are looked for in their ‘standard’ locations (either the top source directory, or in the source directory corresponding to the current Makefile.am, whichever is appropriate). See Section “Finding ‘configure’ Input” in The Autoconf Manual. FIXME: give complete list of things looked for in this directory

AC_CANONICAL_HOST

Automake will ensure that config.guess and config.sub exist. Also, the Makefile variables ‘host_alias’ and ‘host_triplet’ are introduced. See Section “Getting the Canonical System Type” in The Autoconf Manual.

AC_CANONICAL_SYSTEM

This is similar to AC_CANONICAL_HOST, but also defines the Makefile variables ‘build_alias’ and ‘target_alias’. See Section “Getting the Canonical System Type” in The Autoconf Manual.

AC_LIBSOURCE
AC_LIBSOURCES
AC_LIBOBJ

Automake will automatically distribute any file listed in AC_LIBSOURCE or AC_

LIBSOURCES.

Note that the AC_LIBOBJ macro calls AC_LIBSOURCE. So if an Autoconf macro is documented to call AC_LIBOBJ([file]), then file.c will be distributed automatically by Automake. This encompasses many macros like AC_FUNC_

ALLOCA, AC_FUNC_MEMCMP, AC_REPLACE_FUNCS, and others.

By the way, direct assignments to LIBOBJS are no longer supported. You should always use AC_LIBOBJ for this purpose. See Section “AC_LIBOBJ vs. LIBOBJS” in The Autoconf Manual.

AC_PROG_RANLIB

This is required if any libraries are built in the package. See Section “Particular Program Checks” in The Autoconf Manual.

AC_PROG_CXX

This is required if any C++ source is included. See Section “Particular Program Checks” in The Autoconf Manual.

AC_PROG_F77

This is required if any Fortran 77 source is included. This macro is distributed with Autoconf version 2.13 and later. See Section “Particular Program Checks” in The Autoconf Manual.

AC_F77_LIBRARY_LDFLAGS

This is required for programs and shared libraries that are a mixture of languages that include Fortran 77 (see Section 9.10.3 [Mixing Fortran 77 With C and C++], page 35). See Section 5.6 [Autoconf macros supplied with Automake], page 17.

AC_PROG_LIBTOOL

Automake will turn on processing for libtool (see Section “Introduction” in The Libtool Manual).

AC_PROG_YACC

If a Yacc source file is seen, then you must either use this macro or define the variable ‘YACC’ in configure.in. The former is preferred (see Section “Particular Program Checks” in The Autoconf Manual).

AC_PROG_LEX

If a Lex source file is seen, then this macro must be used. See Section “Particular Program Checks” in The Autoconf Manual.

AC_SUBST The first argument is automatically defined as a variable in each generated Makefile.in. See Section “Setting Output Variables” in The Autoconf Manual.

If the Autoconf manual says that a macro calls AC_SUBST for var, or defined the output variable var then var will be defined in each generated Makefile.in. E.g. AC_PATH_XTRA defines X_CFLAGS and X_LIBS, so you can use the variable

          in any Makefile.am if AC_PATH_XTRA is called.
AM_C_PROTOTYPES

This is required when using automatic de-ANSI-fication; see Section 9.13 [ANSI], page 38.

AM_GNU_GETTEXT

This macro is required for packages which use GNU gettext (see Section 11.2 [gettext], page 42). It is distributed with gettext. If Automake sees this macro it ensures that the package meets some of gettext’s requirements.

AM_MAINTAINER_MODE
          This macro adds a ‘--enable-maintainer-mode’ option to configure. If this is

used, automake will cause ‘maintainer-only’ rules to be turned off by default in the generated Makefile.ins. This macro defines the ‘MAINTAINER_MODE’ conditional, which you can use in your own Makefile.am.

5.3 Auto-generating aclocal.m4

Automake includes a number of Autoconf macros which can be used in your package; some of them are actually required by Automake in certain situations. These macros must be defined in your aclocal.m4; otherwise they will not be seen by autoconf.

The aclocal program will automatically generate aclocal.m4 files based on the contents of configure.in. This provides a convenient way to get Automake-provided macros, without having to search around. Also, the aclocal mechanism allows other packages to supply their own macros.

At startup, aclocal scans all the .m4 files it can find, looking for macro definitions (see Section 5.5 [Macro search path], page 15). Then it scans configure.in. Any mention of one of the macros found in the first step causes that macro, and any macros it in turn requires, to be put into aclocal.m4.

The contents of acinclude.m4, if it exists, are also automatically included in aclocal.m4. This is useful for incorporating local macros into configure.

aclocal tries to be smart about looking for new AC_DEFUNs in the files it scans. It also tries to copy the full text of the scanned file into aclocal.m4, including both ‘#’ and ‘dnl’ comments. If you want to make a comment which will be completely ignored by aclocal, use ‘##’ as the comment leader.

5.4 aclocal options

aclocal accepts the following options:

--acdir=dir

Look for the macro files in dir instead of the installation directory. This is typically used for debugging.

--help Print a summary of the command line options and exit.

-I dir Add the directory dir to the list of directories searched for .m4 files.

--output=file

Cause the output to be put into file instead of aclocal.m4.

--print-ac-dir

Prints the name of the directory which aclocal will search to find third-party .m4 files. When this option is given, normal processing is suppressed. This option can be used by a package to determine where to install a macro file.

--verbose

Print the names of the files it examines.

--version

Print the version number of Automake and exit.

5.5 Macro search path

By default, aclocal searches for .m4 files in the following directories, in this order:

acdir-APIVERSION

This is where the .m4 macros distributed with automake itself are stored. APIVERSION depends on the automake release used; for automake 1.6.x, APIVERSION = 1.6.

acdir This directory is intended for third party .m4 files, and is configured when automake itself is built. This is @datadir@/aclocal/, which typically expands to ${prefix}/share/aclocal/. To find the compiled-in value of acdir, use the --print-ac-dir option (see Section 5.4 [aclocal options], page 15).

As an example, suppose that automake-1.6.2 was configured with --prefix=/usr/local. Then, the search path would be:

 1. /usr/local/share/aclocal-1.6/
 2. /usr/local/share/aclocal/

As explained in (see Section 5.4 [aclocal options], page 15), there are several options that can be used to change or extend this search path.

5.5.1 Modifying the macro search path: --acdir

The most obvious option to modify the search path is --acdir=dir, which changes default directory and drops the APIVERSION directory. For example, if one specifies --acdir=/opt/private/, then the search path becomes:

 1. /opt/private/

Note that this option, --acdir, is intended for use by the internal automake test suite only; it is not ordinarily needed by end-users.

5.5.2 Modifying the macro search path: -I dir

Any extra directories specified using -I options (see Section 5.4 [aclocal options], page 15) are prepended to this search list. Thus, aclocal -I /foo -I /bar results in the following search path:

 1. /foo
 2. /bar

3. acdir-APIVERSION

4. acdir

5.5.3 Modifying the macro search path: dirlist

There is a third mechanism for customizing the search path. If a dirlist file exists in acdir, then that file is assumed to contain a list of directories, one per line, to be added to the search list. These directories are searched after all other directories.

For example, suppose acdir/dirlist contains the following:

     /test1
     /test2

and that aclocal was called with the -I /foo -I /bar options. Then, the search path would be

 1. /foo
 2. /bar

3. acdir-APIVERSION

4. acdir

 5. /test1
 6. /test2

If the --acdir=dir option is used, then aclocal will search for the dirlist file in dir. In the --acdir=/opt/private/ example above, aclocal would look for /opt/private/dirlist. Again, however, the --acdir option is intended for use by the internal automake test suite only; --acdir is not ordinarily needed by end-users.

dirlist is useful in the following situation: suppose that automake version 1.6.2 is installed with $prefix=/usr by the system vendor. Thus, the default search directories are

 1. /usr/share/aclocal-1.6/
 2. /usr/share/aclocal/

However, suppose further that many packages have been manually installed on the system, with $prefix=/usr/local, as is typical. In that case, many of these “extra” .m4 files are

in /usr/local/share/aclocal. The only way to force /usr/bin/aclocal to find these

“extra” .m4 files is to always call aclocal -I /usr/local/share/aclocal. This is inconvenient. With dirlist, one may create the file

   /usr/share/aclocal/dirlist

which contains only the single line

/usr/local/share/aclocal

Now, the “default” search path on the affected system is

1. /usr/share/aclocal-1.6/
2. /usr/share/aclocal/
3. /usr/local/share/aclocal/

without the need for -I options; -I options can be reserved for project-specific needs (my-source-dir/m4/), rather than using it to work around local system-dependent tool installation directories.

Similarly, dirlist can be handy if you have installed a local copy Automake on your account and want aclocal to look for macros installed at other places on the system.

5.6 Autoconf macros supplied with Automake

Automake ships with several Autoconf macros that you can use from your configure.in. When you use one of them it will be included by aclocal in aclocal.m4.

5.6.1 Public macros

AM_CONFIG_HEADER

Automake will generate rules to automatically regenerate the config header. This obsolete macro is a synonym of AC_CONFIG_HEADERS today (see Section 5.2 [Optional], page 12).

AM_ENABLE_MULTILIB

This is used when a “multilib” library is being built. The first optional argument is the name of the Makefile being generated; it defaults to ‘Makefile’. The second option argument is used to find the top source directory; it defaults to the empty string (generally this should not be used unless you are familiar with the internals). See Section 18.3 [Multilibs], page 55.

AM_C_PROTOTYPES

Check to see if function prototypes are understood by the compiler. If so, define ‘PROTOTYPES’ and set the output variables ‘U’ and ‘ANSI2KNR’ to the empty string. Otherwise, set ‘U’ to ‘_’ and ‘ANSI2KNR’ to ‘./ansi2knr’. Automake uses these values to implement automatic de-ANSI-fication.

AM_HEADER_TIOCGWINSZ_NEEDS_SYS_IOCTL
If the use of TIOCGWINSZ requires <sys/ioctl.h>, then define GWINSZ_IN_SYS_
IOCTL. Otherwise TIOCGWINSZ can be found in <termios.h>.
AM_INIT_AUTOMAKE([OPTIONS])
AM_INIT_AUTOMAKE(PACKAGE, VERSION, [NO-DEFINE])

Runs many macros required for proper operation of the generated Makefiles. This macro has two forms, the second of which has two required arguments: the package and the version number. This latter form is obsolete because the package and version can be obtained from Autoconf’s AC_INIT macro (which itself has an old and a new form).

If your configure.in has:
     AC_INIT(src/foo.c)

     AM_INIT_AUTOMAKE(mumble, 1.5)

you can modernize it as follow:

     AC_INIT(mumble, 1.5)
     AC_CONFIG_SRCDIR(src/foo.c)
     AM_INIT_AUTOMAKE

Note that if you’re upgrading your configure.in from an earlier version of Automake, it is not always correct to simply move the package and version arguments from AM_INIT_AUTOMAKE directly to AC_INIT, as in the example above. The first argument of AC_INIT is the name of your package (e.g. ‘GNU Automake’), not the tarball name (e.g. ‘automake’) you used to pass to AM_ INIT_AUTOMAKE. Autoconf’s rule to derive a tarball name from the package name should work for most but not all packages. Especially, if your tarball name is not all lower case, you will have to use the four-argument form of AC_INIT (supported in Autoconf versions greater than 2.52g). When AM_INIT_AUTOMAKE is called with a single argument, it is interpreted as a space-separated list of Automake options which should be applied to every Makefile.am in the tree. The effect is as if each option were listed in AUTOMAKE_

OPTIONS.

By default this macro AC_DEFINE’s ‘PACKAGE’ and ‘VERSION’. This can be avoided by passing the ‘no-define’ option, as in:

     AM_INIT_AUTOMAKE([gnits 1.5 no-define dist-bzip2])

or by passing a third non-empty argument to the obsolete form.

AM_PATH_LISPDIR

Searches for the program emacs, and, if found, sets the output variable lispdir to the full path to Emacs’ site-lisp directory. Note that this test assumes the emacs found to be a version that supports Emacs Lisp (such as gnu Emacs or XEmacs). Other emacsen can cause this test to hang (some, like old versions of MicroEmacs, start up in interactive mode, requiring ‘C-x C-c’ to exit, which is hardly obvious for a non-emacs user). In most cases, however, you should be able to use ‘C-c’ to kill the test. In order to avoid problems, you can set EMACS to “no” in the environment, or use the ‘--with-lispdir’ option to configure to explictly set the correct path (if you’re sure you have an emacs that supports Emacs Lisp.

AM_PROG_AS

Use this macro when you have assembly code in your project. This will choose the assembler for you (by default the C compiler) and set CCAS, and will also set CCASFLAGS if required.

AM_PROG_CC_C_O

This is like AC_PROG_CC_C_O, but it generates its results in the manner required by automake. You must use this instead of AC_PROG_CC_C_O when you need this functionality.

AM_PROG_CC_STDC

If the C compiler is not in ANSI C mode by default, try to add an option to output variable CC to make it so. This macro tries various options that select ANSI C on some system or another. It considers the compiler to be in ANSI C mode if it handles function prototypes correctly. If you use this macro, you should check after calling it whether the C compiler has been set to accept ANSI C; if not, the shell variable am_cv_prog_cc_stdc is set to ‘no’. If you wrote your source code in ANSI C, you can make an un-ANSIfied copy of it by using the ansi2knr option (see Section 9.13 [ANSI], page 38).

AM_PROG_LEX

Like AC_PROG_LEX (see Section “Particular Program Checks” in The Autoconf Manual), but uses the missing script on systems that do not have lex. ‘HP-UX 10’ is one such system.

AM_PROG_GCJ

This macro finds the gcj program or causes an error. It sets ‘GCJ’ and ‘GCJFLAGS’. gcj is the Java front-end to the GNU Compiler Collection.

AM_SYS_POSIX_TERMIOS

Check to see if POSIX termios headers and functions are available on the system. If so, set the shell variable am_cv_sys_posix_termios to ‘yes’. If not, set the variable to ‘no’.

AM_WITH_DMALLOC
Add support for the dmalloc (ftp://ftp.letters.com/src/dmalloc/

dmalloc.tar.gz) package. If the user configures with ‘--with-dmalloc’, then

define WITH_DMALLOC and add ‘-ldmalloc’ to LIBS.
AM_WITH_REGEX

Adds ‘--with-regex’ to the configure command line. If specified (the default), then the ‘regex’ regular expression library is used, regex.o is put into ‘LIBOBJS’, and ‘WITH_REGEX’ is defined. If ‘--without-regex’ is given, then the ‘rx’ regular expression library is used, and rx.o is put into ‘LIBOBJS’.

5.6.2 Private macros

The following macros are private macros you should not call directly. They are called by the other public macros when appropriate. Do not rely on them, as they might be changed in a future version. Consider them as implementation details; or better, do not consider them at all: skip this section!

_AM_DEPENDENCIES
AM_SET_DEPDIR
AM_DEP_TRACK
AM_OUTPUT_DEPENDENCY_COMMANDS

These macros are used to implement automake’s automatic dependency tracking scheme. They are called automatically by automake when required, and there should be no need to invoke them manually.

AM_MAKE_INCLUDE

This macro is used to discover how the user’s make handles include statements. This macro is automatically invoked when needed; there should be no need to invoke it manually.

AM_PROG_INSTALL_STRIP

This is used to find a version of install which can be used to strip a program at installation time. This macro is automatically included when required.

AM_SANITY_CHECK

This checks to make sure that a file created in the build directory is newer than a file in the source directory. This can fail on systems where the clock is set incorrectly. This macro is automatically run from AM_INIT_AUTOMAKE.

5.7 Writing your own aclocal macros

The aclocal program doesn’t have any built-in knowledge of any macros, so it is easy to extend it with your own macros.

This is mostly used for libraries which want to supply their own Autoconf macros for use by other programs. For instance the gettext library supplies a macro AM_GNU_GETTEXT which should be used by any package using gettext. When the library is installed, it installs this macro so that aclocal will find it.

A file of macros should be a series of AC_DEFUN’s. The aclocal programs also understands AC_REQUIRE, so it is safe to put each macro in a separate file. See Section “Prerequisite Macros” in The Autoconf Manual, and Section “Macro Definitions” in The Autoconf Manual.

A macro file’s name should end in .m4. Such files should be installed in `aclocal

--print-ac-dir` (which usually happens to be $(datadir)/aclocal).

6 The top-level Makefile.am

6.1 Recursing subdirectories

In packages with subdirectories, the top level Makefile.am must tell Automake which subdirectories are to be built. This is done via the SUBDIRS variable.

The SUBDIRS variable holds a list of subdirectories in which building of various sorts can occur. Many targets (e.g. all) in the generated Makefile will run both locally and in all specified subdirectories. Note that the directories listed in SUBDIRS are not required to contain Makefile.ams; only Makefiles (after configuration). This allows inclusion of libraries from packages which do not use Automake (such as gettext).

In packages that use subdirectories, the top-level Makefile.am is often very short. For instance, here is the Makefile.am from the GNU Hello distribution:

     EXTRA_DIST = BUGS ChangeLog.O README-alpha
     SUBDIRS = doc intl po src tests

When Automake invokes make in a subdirectory, it uses the value of the MAKE variable. It passes the value of the variable AM_MAKEFLAGS to the make invocation; this can be set in Makefile.am if there are flags you must always pass to make.

The directories mentioned in SUBDIRS must be direct children of the current directory. For instance, you cannot put ‘src/subdir’ into SUBDIRS. Instead you should put SUBDIRS = subdir into src/Makefile.am. Automake can be used to construct packages of arbitrary depth this way.

By default, Automake generates Makefiles which work depth-first (‘postfix’). However, it is possible to change this ordering. You can do this by putting ‘.’ into SUBDIRS. For instance, putting ‘.’ first will cause a ‘prefix’ ordering of directories. All ‘clean’ targets are run in reverse order of build targets.

6.2 Conditional subdirectories

It is possible to define the SUBDIRS variable conditionally if, like in the case of GNU Inetutils, you want to only build a subset of the entire package.

To illustrate how this works, let’s assume we have two directories src/ and opt/. src/ should always be built, but we want to decide in ./configure whether opt/ will be built or not. (For this example we will assume that opt/ should be built when the variable

$want_opt was set to yes.)

Running make should thus recurse into src/ always, and then maybe in opt/.

However make dist should always recurse into both src/ and opt/. Because opt/ should be distributed even if it is not needed in the current configuration. This means opt/Makefile should be created unconditionally.^{3}

There are two ways to setup a project like this. You can use Automake conditionals (see Chapter 20 [Conditionals], page 56) or use Autoconf AC_SUBST variables (see Section “Setting Output Variables” in The Autoconf Manual). Using Automake conditionals is the preferred solution.

6.2.1 Conditional subdirectories with AM_CONDITIONAL

configure should output the Makefile for each directory and define a condition into which opt/ should be built.

     ...
     AM_CONDITIONAL([COND_OPT], [test "$want_opt" = yes])
     AC_CONFIG_FILES([Makefile src/Makefile opt/Makefile])
     ...

Then SUBDIRS can be defined in the top-level Makefile.am as follows.

     if COND_OPT
       MAYBE_OPT = opt
     endif
     SUBDIRS = src $(MAYBE_OPT)

As you can see, running make will rightly recurse into src/ and maybe opt/.

As you can’t see, running make dist will recurse into both src/ and opt/ directories because make dist, unlike make all, doesn’t use the SUBDIRS variable. It uses the DIST_ SUBDIRS variable.

In this case Automake will define DIST_SUBDIRS = src opt automatically because it knows that MAYBE_OPT can contain opt in some condition.

3 Don’t try seeking a solution where opt/Makefile is created conditionally, this is a lot trickier than the

solutions presented here.

6.2.2 Conditional subdirectories with AC_SUBST

Another idea is to define MAYBE_OPT from ./configure using AC_SUBST:

     ...
     if test "$want_opt" = yes; then
       MAYBE_OPT=opt
     else
       MAYBE_OPT=
     fi
     AC_SUBST([MAYBE_OPT])
     AC_CONFIG_FILES([Makefile src/Makefile opt/Makefile])
     ...

In this case the top-level Makefile.am should look as follows.

     SUBDIRS = src $(MAYBE_OPT)
     DIST_SUBDIRS = src opt

The drawback is that since Automake cannot guess what the possible values of MAYBE_ OPT are, it is necessary to define DIST_SUBDIRS.

6.2.3 How DIST_SUBDIRS is used

As shown in the above examples, DIST_SUBDIRS is used by targets that need to recurse in all directories, even those which have been conditionally left out of the build.

   Precisely, DIST_SUBDIRS is used by make dist, make distclean, and make

maintainer-clean. All other recursive targets use SUBDIRS.

Automake will define DIST_SUBDIRS automatically from the possibles values of SUBDIRS in all conditions.

If SUBDIRS contains AC_SUBST variables, DIST_SUBDIRS will not be defined correctly because Automake doesn’t know the possible values of these variables. In this case DIST_ SUBDIRS needs to be defined manually.

7 An Alternative Approach to Subdirectories

If you’ve ever read Peter Miller’s excellent paper, Recursive Make Considered Harmful

(http://www.pcug.org.au/~millerp/rmch/recu-make-cons-harm.html), the preceding

section on the use of subdirectories will probably come as unwelcome advice. For those who haven’t read the paper, Miller’s main thesis is that recursive make invocations are both slow and error-prone.

Automake provides sufficient cross-directory support^{4} to enable you to write a single Makefile.am for a complex multi-directory package.

By default an installable file specified in a subdirectory will have its directory name stripped before installation. For instance, in this example, the header file will be installed

as $(includedir)/stdio.h:
     include_HEADERS = inc/stdio.h

4 We believe. This work is new and there are probably warts. See Chapter 1 [Introduction], page 1, for

information on reporting bugs.

However, the ‘nobase_’ prefix can be used to circumvent this path stripping. In this example, the header file will be installed as $(includedir)/sys/types.h:

     nobase_include_HEADERS = sys/types.h

‘nobase_’ should be specified first when used in conjonction with either ‘dist_’ or ‘nodist_’ (see Chapter 15 [Dist], page 48). For instance:

     nobase_dist_pkgdata_DATA = images/vortex.pgm

8 Rebuilding Makefiles

Automake generates rules to automatically rebuild Makefiles, configure, and other derived files like Makefile.in.

If you are using AM_MAINTAINER_MODE in configure.in, then these automatic rebuilding rules are only enabled in maintainer mode.

Sometimes you need to run aclocal with an argument like -I to tell it where to find .m4 files. Since sometimes make will automatically run aclocal, you need a way to specify these arguments. You can do this by defining ACLOCAL_AMFLAGS; this holds arguments which are passed verbatim to aclocal. This variable is only useful in the top-level Makefile.am.

9 Building Programs and Libraries

A large part of Automake’s functionality is dedicated to making it easy to build programs and libraries.

9.1 Building a program

In order to build a program, you need to tell Automake which sources are part of it, and which libraries it should be linked with.

This section also covers conditional compilation of sources or programs. Most of the comments about these also apply to libraries (see Section 9.2 [A Library], page 26) and Libtool libraries (see Section 9.3 [A Shared Library], page 27).

9.1.1 Defining program sources

In a directory containing source that gets built into a program (as opposed to a library or a script), the ‘PROGRAMS’ primary is used. Programs can be installed in bindir, sbindir, libexecdir, pkglibdir, or not at all (‘noinst’). They can also be built only for make check, in which case the prefix is ‘check’.

For instance:

     bin_PROGRAMS = hello

In this simple case, the resulting Makefile.in will contain code to generate a program named hello.

Associated with each program are several assisting variables which are named after the program. These variables are all optional, and have reasonable defaults. Each variable, its use, and default is spelled out below; we use the “hello” example throughout.

The variable hello_SOURCES is used to specify which source files get built into an executable:

     hello_SOURCES = hello.c version.c getopt.c getopt1.c getopt.h system.h

This causes each mentioned ‘.c’ file to be compiled into the corresponding ‘.o’. Then all are linked to produce hello.

If ‘hello_SOURCES’ is not specified, then it defaults to the single file hello.c; that is, the default is to compile a single C file whose base name is the name of the program itself. (This is a terrible default but we are stuck with it for historical reasons.)

Multiple programs can be built in a single directory. Multiple programs can share a single source file, which must be listed in each ‘_SOURCES’ definition.

Header files listed in a ‘_SOURCES’ definition will be included in the distribution but otherwise ignored. In case it isn’t obvious, you should not include the header file generated by configure in a ‘_SOURCES’ variable; this file should not be distributed. Lex (‘.l’) and Yacc (‘.y’) files can also be listed; see Section 9.7 [Yacc and Lex], page 31.

9.1.2 Linking the program

If you need to link against libraries that are not found by configure, you can use LDADD to do so. This variable is used to specify additional objects or libraries to link with; it is inappropriate for specifying specific linker flags, you should use AM_LDFLAGS for this purpose.

Sometimes, multiple programs are built in one directory but do not share the same linktime requirements. In this case, you can use the ‘prog_LDADD’ variable (where prog is the name of the program as it appears in some ‘_PROGRAMS’ variable, and usually written in lowercase) to override the global LDADD. If this variable exists for a given program, then that program is not linked using LDADD.

For instance, in GNU cpio, pax, cpio and mt are linked against the library libcpio.a. However, rmt is built in the same directory, and has no such link requirement. Also, mt and rmt are only built on certain architectures. Here is what cpio’s src/Makefile.am looks like (abridged):

     bin_PROGRAMS = cpio pax @MT@
     libexec_PROGRAMS = @RMT@
     EXTRA_PROGRAMS = mt rmt
     LDADD = ../lib/libcpio.a @INTLLIBS@
     rmt_LDADD =
     cpio_SOURCES = ...
     pax_SOURCES = ...
     mt_SOURCES = ...
     rmt_SOURCES = ...

‘prog_LDADD’ is inappropriate for passing program-specific linker flags (except for ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’). So, use the ‘prog_LDFLAGS’ variable for this purpose.

It is also occasionally useful to have a program depend on some other target which is not actually part of that program. This can be done using the ‘prog_DEPENDENCIES’ variable.

Each program depends on the contents of such a variable, but no further interpretation is done.

If ‘prog_DEPENDENCIES’ is not supplied, it is computed by Automake. The automatically-assigned value is the contents of ‘prog_LDADD’, with most configure substitutions, ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’ options removed. The configure substitutions that are left in are only ‘@LIBOBJS@’ and ‘@ALLOCA@’; these are left because it is known that they will not cause an invalid value for ‘prog_DEPENDENCIES’ to be generated.

9.1.3 Conditional compilation of sources

You can’t put a configure substitution (e.g., ‘@FOO@’) into a ‘_SOURCES’ variable. The reason for this is a bit hard to explain, but suffice to say that it simply won’t work. Automake will give an error if you try to do this.

Fortunatly there are two other ways to achieve the same result. One is to use configure substitutions in _LDADD variables, the other is to use an Automake conditional.

9.1.3.1 Conditional compilation using _LDADD substitutions

Automake must know all the source files that could possibly go into a program, even if not all the files are built in every circumstance. Any files which are only conditionally built should be listed in the appropriate ‘EXTRA_’ variable. For instance, if hello-linux.c or hello-generic.c were conditionally included in hello, the Makefile.am would contain:

     bin_PROGRAMS = hello
     hello_SOURCES = hello-common.c
     EXTRA_hello_SOURCES = hello-linux.c hello-generic.c
     hello_LDADD = @HELLO_SYSTEM@
     hello_DEPENDENCIES = @HELLO_SYSTEM@

You can then setup the @HELLO_SYSTEM@ substitution from configure.in:

     ...
     case $host in
       *linux*) HELLO_SYSTEM='hello-linux.$(OBJEXT)' ;;
       *)       HELLO_SYSTEM='hello-generic.$(OBJEXT)' ;;
     esac
     AC_SUBST([HELLO_SYSTEM])
     ...
   In this case, HELLO_SYSTEM should be replaced by hello-linux.o or hello-bsd.o, and

added to hello_DEPENDENCIES and hello_LDADD in order to be built and linked in.

9.1.3.2 Conditional compilation using Automake conditionals

An often simpler way to compile source files conditionally is to use Automake conditionals. For instance, you could use this Makefile.am construct to build the same hello example:

     bin_PROGRAMS = hello
     if LINUX
     hello_SOURCES = hello-linux.c hello-common.c
     else
     hello_SOURCES = hello-generic.c hello-common.c

     endif

In this case, your configure.in should setup the LINUX conditional using AM_CONDITIONAL (see Chapter 20 [Conditionals], page 56).

When using conditionals like this you don’t need to use the ‘EXTRA_’ variable, because Automake will examine the contents of each variable to construct the complete list of source files.

If your program uses a lot of files, you will probably prefer a conditional +=.

     bin_PROGRAMS = hello
     hello_SOURCES = hello-common.c
     if LINUX
     hello_cond += hello-linux.c
     else
     hello_cond += hello-generic.c
     endif

9.1.4 Conditional compilation of programs

Sometimes it is useful to determine the programs that are to be built at configure time. For instance, GNU cpio only builds mt and rmt under special circumstances.

In this case, you must notify Automake of all the programs that can possibly be built, but at the same time cause the generated Makefile.in to use the programs specified by configure. This is done by having configure substitute values into each ‘_PROGRAMS’ definition, while listing all optionally built programs in EXTRA_PROGRAMS.

Of course you can use Automake conditionals to determine the programs to be built.

9.2 Building a library

Building a library is much like building a program. In this case, the name of the primary is ‘LIBRARIES’. Libraries can be installed in libdir or pkglibdir.

See Section 9.3 [A Shared Library], page 27, for information on how to build shared libraries using Libtool and the ‘LTLIBRARIES’ primary.

Each ‘_LIBRARIES’ variable is a list of the libraries to be built. For instance to create a library named libcpio.a, but not install it, you would write:

     noinst_LIBRARIES = libcpio.a

The sources that go into a library are determined exactly as they are for programs, via the ‘_SOURCES’ variables. Note that the library name is canonicalized (see Section 2.4 [Canonicalization], page 4), so the ‘_SOURCES’ variable corresponding to liblob.a is

‘liblob_a_SOURCES’, not ‘liblob.a_SOURCES’.

Extra objects can be added to a library using the ‘library_LIBADD’ variable. This should be used for objects determined by configure. Again from cpio:

     libcpio_a_LIBADD = @LIBOBJS@ @ALLOCA@

In addition, sources for extra objects that will not exist until configure-time must be added to the BUILT_SOURCES variable (see Section 10.4 [Sources], page 41).

9.3 Building a Shared Library

Building shared libraries is a relatively complex matter. For this reason, GNU Libtool (see Section “Introduction” in The Libtool Manual) was created to help build shared libraries in a platform-independent way.

Automake uses Libtool to build libraries declared with the ‘LTLIBRARIES’ primary. Each ‘_LTLIBRARIES’ variable is a list of shared libraries to build. For instance, to create a library named libgettext.a and its corresponding shared libraries, and install them in ‘libdir’, write:

     lib_LTLIBRARIES = libgettext.la

Note that shared libraries must be installed in order to work properly, so check_ LTLIBRARIES is not allowed. However, noinst_LTLIBRARIES is allowed. This feature should be used for libtool “convenience libraries”.

For each library, the ‘library_LIBADD’ variable contains the names of extra libtool objects (.lo files) to add to the shared library. The ‘library_LDFLAGS’ variable contains any additional libtool flags, such as ‘-version-info’ or ‘-static’.

Where an ordinary library might include @LIBOBJS@, a libtool library must use @LTLIBOBJS@. This is required because the object files that libtool operates on do not necessarily end in .o. The libtool manual contains more details on this topic.

For libraries installed in some directory, Automake will automatically supply the appropriate ‘-rpath’ option. However, for libraries determined at configure time (and thus mentioned in EXTRA_LTLIBRARIES), Automake does not know the eventual installation directory; for such libraries you must add the ‘-rpath’ option to the appropriate ‘_LDFLAGS’ variable by hand.

Ordinarily, Automake requires that a shared library’s name start with ‘lib’. However, if you are building a dynamically loadable module then you might wish to use a "nonstandard" name. In this case, put -module into the ‘_LDFLAGS’ variable.

See Section “The Libtool Manual” in The Libtool Manual, for more information.

9.4 Program and Library Variables

Associated with each program are a collection of variables which can be used to modify how that program is built. There is a similar list of such variables for each library. The canonical name of the program (or library) is used as a base for naming these variables.

In the list below, we use the name “maude” to refer to the program or library. In your Makefile.am you would replace this with the canonical name of your program. This list also refers to “maude” as a program, but in general the same rules apply for both static and dynamic libraries; the documentation below notes situations where programs and libraries differ.

‘maude_SOURCES’

This variable, if it exists, lists all the source files which are compiled to build the program. These files are added to the distribution by default. When building the program, Automake will cause each source file to be compiled to a single .o file (or .lo when using libtool). Normally these object files are named after the source file, but other factors can change this. If a file in the ‘_SOURCES’ variable has an unrecognized extension, Automake will do one of two things with it. If a suffix rule exists for turning files with the unrecognized extension into .o files, then automake will treat this file as it will any other source file (see Section 9.12 [Support for Other Languages], page 38). Otherwise, the file will be ignored as though it were a header file.

The prefixes ‘dist_’ and ‘nodist_’ can be used to control whether files listed in a ‘_SOURCES’ variable are distributed. ‘dist_’ is redundant, as sources are distributed by default, but it can be specified for clarity if desired.

It is possible to have both ‘dist_’ and ‘nodist_’ variants of a given ‘_SOURCES’ variable at once; this lets you easily distribute some files and not others, for instance:

     nodist_maude_SOURCES = nodist.c
     dist_maude_SOURCES = dist-me.c

By default the output file (on Unix systems, the .o file) will be put into the current build directory. However, if the option subdir-objects is in effect in the current directory then the .o file will be put into the subdirectory named after the source file. For instance, with subdir-objects enabled, sub/dir/file.c will be compiled to sub/dir/file.o. Some people prefer this mode of operation. You can specify subdir-objects in AUTOMAKE_OPTIONS (see Chapter 17 [Options], page 52).

‘EXTRA_maude_SOURCES’

Automake needs to know the list of files you intend to compile statically. For one thing, this is the only way Automake has of knowing what sort of language support a given Makefile.in requires.^{5} This means that, for example, you can’t put a configure substitution like ‘@my_sources@’ into a ‘_SOURCES’ variable. If you intend to conditionally compile source files and use configure to substitute the appropriate object names into, e.g., ‘_LDADD’ (see below), then you should list the corresponding source files in the ‘EXTRA_’ variable.

This variable also supports ‘dist_’ and ‘nodist_’ prefixes, e.g.,

‘nodist_EXTRA_maude_SOURCES’.
‘maude_AR’

A static library is created by default by invoking $(AR) cru followed by the name of the library and then the objects being put into the library. You can override this by setting the ‘_AR’ variable. This is usually used with C++; some C++ compilers require a special invocation in order to instantiate all the templates which should go into a library. For instance, the SGI C++ compiler likes this variable set like so:

     libmaude_a_AR = $(CXX) -ar -o
‘maude_LIBADD’

Extra objects can be added to a library using the ‘_LIBADD’ variable. This should be used for objects determined by configure.

5 There are other, more obscure reasons reasons for this limitation as well.

‘maude_LDADD’

Extra objects can be added to a shared library or a program by listing them in the ‘_LDADD’ variable. This should be used for objects determined by

configure.

‘_LDADD’ and ‘_LIBADD’ are inappropriate for passing program-specific linker flags (except for ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’). Use the ‘_LDFLAGS’ variable for this purpose. For instance, if your configure.in uses AC_PATH_XTRA, you could link your program against the X libraries like so:

     maude_LDADD = $(X_PRE_LIBS) $(X_LIBS) $(X_EXTRA_LIBS)
‘maude_LDFLAGS’

This variable is used to pass extra flags to the link step of a program or a shared library.

‘maude_LINK’

You can override the linker on a per-program basis. By default the linker is chosen according to the languages used by the program. For instance, a program that includes C++ source code would use the C++ compiler to link. The ‘_LINK’ variable must hold the name of a command which can be passed all the .o file names as arguments. Note that the name of the underlying program is not passed to ‘_LINK’; typically one uses ‘$@’:

     maude_LINK = $(CCLD) -magic -o $@
‘maude_CCASFLAGS’
‘maude_CFLAGS’
‘maude_CPPFLAGS’
‘maude_CXXFLAGS’
‘maude_FFLAGS’
‘maude_GCJFLAGS’
‘maude_LFLAGS’
‘maude_OBJCFLAGS’
‘maude_RFLAGS’
‘maude_YFLAGS’

Automake allows you to set compilation flags on a per-program (or per-library) basis. A single source file can be included in several programs, and it will potentially be compiled with different flags for each program. This works for any language directly supported by Automake. The flags are ‘_CCASFLAGS’, ‘_CFLAGS’,

‘_CPPFLAGS’, ‘_CXXFLAGS’, ‘_FFLAGS’, ‘_GCJFLAGS’, ‘_LFLAGS’, ‘_OBJCFLAGS’,
‘_RFLAGS’, and ‘_YFLAGS’.

When using a per-program compilation flag, Automake will choose a different name for the intermediate object files. Ordinarily a file like sample.c will be compiled to produce sample.o. However, if the program’s ‘_CFLAGS’ variable is set, then the object file will be named, for instance, maude-sample.o. In compilations with per-program flags, the ordinary ‘AM_’ form of the flags variable is not automatically included in the compilation (however, the user form of the variable is included). So for instance, if you want the hypothetical maude compilations to also use the value of ‘AM_CFLAGS’, you would need to write:

     maude_CFLAGS = ... your flags ... $(AM_CFLAGS)
‘maude_DEPENDENCIES’

It is also occasionally useful to have a program depend on some other target which is not actually part of that program. This can be done using the ‘_DEPENDENCIES’ variable. Each program depends on the contents of such a variable, but no further interpretation is done. If ‘_DEPENDENCIES’ is not supplied, it is computed by Automake. The automatically-assigned value is the contents of ‘_LDADD’ or ‘_LIBADD’, with most configure substitutions, ‘-l’, ‘-L’, ‘-dlopen’ and ‘-dlpreopen’ options removed. The configure substitutions that are left in are only ‘@LIBOBJS@’ and ‘@ALLOCA@’; these are left because it is known that they will not cause an invalid value for ‘_DEPENDENCIES’ to be generated.

‘maude_SHORTNAME’

On some platforms the allowable file names are very short. In order to support these systems and per-program compilation flags at the same time, Automake allows you to set a “short name” which will influence how intermediate object files are named. For instance, if you set ‘maude_SHORTNAME’ to ‘m’, then in the above per-program compilation flag example the object file would be named m-sample.o rather than maude-sample.o. This facility is rarely needed in practice, and we recommend avoiding it until you find it is required.

9.5 Special handling for LIBOBJS and ALLOCA

Automake explicitly recognizes the use of @LIBOBJS@ and @ALLOCA@, and uses this information, plus the list of LIBOBJS files derived from configure.in to automatically include the appropriate source files in the distribution (see Chapter 15 [Dist], page 48). These source files are also automatically handled in the dependency-tracking scheme; see See Section 9.14 [Dependencies], page 39.

@LIBOBJS@ and @ALLOCA@ are specially recognized in any ‘_LDADD’ or ‘_LIBADD’ variable.

9.6 Variables used when building a program

Occasionally it is useful to know which Makefile variables Automake uses for compilations; for instance you might need to do your own compilation in some special cases.

Some variables are inherited from Autoconf; these are CC, CFLAGS, CPPFLAGS, DEFS,

LDFLAGS, and LIBS.

There are some additional variables which Automake itself defines:

AM_CPPFLAGS

The contents of this variable are passed to every compilation which invokes the C preprocessor; it is a list of arguments to the preprocessor. For instance, ‘-I’ and ‘-D’ options should be listed here. Automake already provides some ‘-I’ options automatically. In particular it generates ‘-I$(srcdir)’, ‘-I.’, and a ‘-I’ pointing to the directory holding

          config.h (if you’ve used AC_CONFIG_HEADERS or AM_CONFIG_HEADER). You

can disable the default ‘-I’ options using the ‘nostdinc’ option.

AM_CPPFLAGS is ignored in preference to a per-executable (or per-library) _ CPPFLAGS variable if it is defined.

INCLUDES This does the same job as ‘AM_CPPFLAGS’. It is an older name for the same functionality. This variable is deprecated; we suggest using ‘AM_CPPFLAGS’ instead.

AM_CFLAGS

This is the variable which the Makefile.am author can use to pass in additional C compiler flags. It is more fully documented elsewhere. In some situations, this is not used, in preference to the per-executable (or per-library) _CFLAGS.

COMPILE This is the command used to actually compile a C source file. The filename is appended to form the complete command line.

AM_LDFLAGS

This is the variable which the Makefile.am author can use to pass in additional linker flags. In some situations, this is not used, in preference to the perexecutable (or per-library) _LDFLAGS.

LINK This is the command used to actually link a C program. It already includes ‘-o $@’ and the usual variable references (for instance, CFLAGS); it takes as “arguments” the names of the object files and libraries to link in.

9.7 Yacc and Lex support

Automake has somewhat idiosyncratic support for Yacc and Lex.

Automake assumes that the .c file generated by yacc (or lex) should be named using the basename of the input file. That is, for a yacc source file foo.y, Automake will cause the intermediate file to be named foo.c (as opposed to y.tab.c, which is more traditional).

The extension of a yacc source file is used to determine the extension of the resulting ‘C’ or ‘C++’ file. Files with the extension ‘.y’ will be turned into ‘.c’ files; likewise, ‘.yy’ will become ‘.cc’; ‘.y++’, ‘c++’; and ‘.yxx’, ‘.cxx’.

Likewise, lex source files can be used to generate ‘C’ or ‘C++’; the extensions ‘.l’, ‘.ll’, ‘.l++’, and ‘.lxx’ are recognized.

You should never explicitly mention the intermediate (‘C’ or ‘C++’) file in any ‘SOURCES’ variable; only list the source file.

The intermediate files generated by yacc (or lex) will be included in any distribution that is made. That way the user doesn’t need to have yacc or lex.

If a yacc source file is seen, then your configure.in must define the variable ‘YACC’. This is most easily done by invoking the macro ‘AC_PROG_YACC’ (see Section “Particular Program Checks” in The Autoconf Manual).

When yacc is invoked, it is passed ‘YFLAGS’ and ‘AM_YFLAGS’. The former is a user variable and the latter is intended for the Makefile.am author.

Similarly, if a lex source file is seen, then your configure.in must define the variable ‘LEX’. You can use ‘AC_PROG_LEX’ to do this (see Section “Particular Program Checks” in The Autoconf Manual), but using AM_PROG_LEX macro (see Section 5.6 [Macros], page 17) is recommended.

When lex is invoked, it is passed ‘LFLAGS’ and ‘AM_LFLAGS’. The former is a user variable and the latter is intended for the Makefile.am author.

Automake makes it possible to include multiple yacc (or lex) source files in a single program. When there is more than one distinct yacc (or lex) source file in a directory, Automake uses a small program called ylwrap to run yacc (or lex) in a subdirectory. This is necessary because yacc’s output filename is fixed, and a parallel make could conceivably invoke more than one instance of yacc simultaneously. The ylwrap program is distributed with Automake. It should appear in the directory specified by ‘AC_CONFIG_AUX_DIR’ (see Section “Finding ‘configure’ Input” in The Autoconf Manual), or the current directory if that macro is not used in configure.in.

For yacc, simply managing locking is insufficient. The output of yacc always uses the same symbol names internally, so it isn’t possible to link two yacc parsers into the same executable.

We recommend using the following renaming hack used in gdb:

#define yymaxdepth c_maxdepth
#define yyparse c_parse
#define yylex c_lex
#define yyerror c_error
#define yylval c_lval
#define yychar c_char
#define yydebug c_debug
#define yypact c_pact
#define yyr1 c_r1
#define yyr2 c_r2
#define yydef c_def
#define yychk c_chk
#define yypgo c_pgo
#define yyact c_act
#define yyexca c_exca
#define yyerrflag c_errflag
#define yynerrs c_nerrs
#define yyps c_ps
#define yypv c_pv
#define yys c_s
#define yy_yys c_yys
#define yystate c_state
#define yytmp c_tmp
#define yyv c_v
#define yy_yyv c_yyv
#define yyval c_val
#define yylloc c_lloc
#define yyreds c_reds
#define yytoks c_toks
#define yylhs c_yylhs

     #define yylen c_yylen
     #define yydefred c_yydefred
     #define yydgoto c_yydgoto
     #define yysindex c_yysindex
     #define yyrindex c_yyrindex
     #define yygindex c_yygindex
     #define yytable  c_yytable
     #define yycheck  c_yycheck
     #define yyname   c_yyname
     #define yyrule   c_yyrule

For each define, replace the ‘c_’ prefix with whatever you like. These defines work for bison, byacc, and traditional yaccs. If you find a parser generator that uses a symbol not covered here, please report the new name so it can be added to the list.

9.8 C++ Support

Automake includes full support for C++.

Any package including C++ code must define the output variable ‘CXX’ in configure.in; the simplest way to do this is to use the AC_PROG_CXX macro (see Section “Particular Program Checks” in The Autoconf Manual).

A few additional variables are defined when a C++ source file is seen:

CXX The name of the C++ compiler.

CXXFLAGS Any flags to pass to the C++ compiler.

AM_CXXFLAGS

The maintainer’s variant of CXXFLAGS.

CXXCOMPILE

The command used to actually compile a C++ source file. The file name is appended to form the complete command line.

CXXLINK The command used to actually link a C++ program.

9.9 Assembly Support

Automake includes some support for assembly code.

The variable CCAS holds the name of the compiler used to build assembly code. This compiler must work a bit like a C compiler; in particular it must accept ‘-c’ and ‘-o’. The value of CCASFLAGS is passed to the compilation.

You are required to set CCAS and CCASFLAGS via configure.in. The autoconf macro AM_PROG_AS will do this for you. Unless they are already set, it simply sets CCAS to the C compiler and CCASFLAGS to the C compiler flags.

Only the suffixes ‘.s’ and ‘.S’ are recognized by automake as being files containing assembly code.

9.10 Fortran 77 Support

Automake includes full support for Fortran 77.

Any package including Fortran 77 code must define the output variable ‘F77’ in configure.in; the simplest way to do this is to use the AC_PROG_F77 macro (see Section “Particular Program Checks” in The Autoconf Manual). See Section 9.10.4 [Fortran 77 and Autoconf], page 37.

A few additional variables are defined when a Fortran 77 source file is seen:

F77 The name of the Fortran 77 compiler.

FFLAGS Any flags to pass to the Fortran 77 compiler.

AM_FFLAGS

The maintainer’s variant of FFLAGS.

RFLAGS Any flags to pass to the Ratfor compiler.

AM_RFLAGS

The maintainer’s variant of RFLAGS.

F77COMPILE

The command used to actually compile a Fortran 77 source file. The file name is appended to form the complete command line.

FLINK The command used to actually link a pure Fortran 77 program or shared library.

Automake can handle preprocessing Fortran 77 and Ratfor source files in addition to compiling them^{6}. Automake also contains some support for creating programs and shared libraries that are a mixture of Fortran 77 and other languages (see Section 9.10.3 [Mixing Fortran 77 With C and C++], page 35).

These issues are covered in the following sections.

9.10.1 Preprocessing Fortran 77

N.f is made automatically from N.F or N.r. This rule runs just the preprocessor to convert a preprocessable Fortran 77 or Ratfor source file into a strict Fortran 77 source file. The precise command used is as follows:

.F        $(F77) -F $(DEFS) $(INCLUDES) $(AM_CPPFLAGS) $(CPPFLAGS) $(AM_
          FFLAGS) $(FFLAGS)
.r        $(F77) -F $(AM_FFLAGS) $(FFLAGS) $(AM_RFLAGS) $(RFLAGS)

9.10.2 Compiling Fortran 77 Files

N.o is made automatically from N.f, N.F or N.r by running the Fortran 77 compiler. The precise command used is as follows:

.f        $(F77) -c $(AM_FFLAGS) $(FFLAGS)
.F        $(F77) -c $(DEFS) $(INCLUDES) $(AM_CPPFLAGS) $(CPPFLAGS) $(AM_
          FFLAGS) $(FFLAGS)
.r        $(F77) -c $(AM_FFLAGS) $(FFLAGS) $(AM_RFLAGS) $(RFLAGS)

6 Much, if not most, of the information in the following sections pertaining to preprocessing Fortran 77

programs was taken almost verbatim from Section “Catalogue of Rules” in The GNU Make Manual.

9.10.3 Mixing Fortran 77 With C and C++

Automake currently provides limited support for creating programs and shared libraries that are a mixture of Fortran 77 and C and/or C++. However, there are many other issues related to mixing Fortran 77 with other languages that are not (currently) handled by Automake, but that are handled by other packages^{7}.

7 For example, the cfortran package (http://www-zeus.desy.de/~burow/cfortran/) addresses all of these

inter-language issues, and runs under nearly all Fortran 77, C and C++ compilers on nearly all platforms.

However, cfortran is not yet Free Software, but it will be in the next major release.

Automake can help in two ways:

1. Automatic selection of the linker depending on which combinations of source code.

2. Automatic selection of the appropriate linker flags (e.g. ‘-L’ and ‘-l’) to pass to the automatically selected linker in order to link in the appropriate Fortran 77 intrinsic and run-time libraries. These extra Fortran 77 linker flags are supplied in the output variable FLIBS by the AC_F77_LIBRARY_LDFLAGS Autoconf macro supplied with newer versions of Autoconf (Autoconf version 2.13 and later). See Section “Fortran 77 Compiler Characteristics” in The Autoconf .

If Automake detects that a program or shared library (as mentioned in some _PROGRAMS or _LTLIBRARIES primary) contains source code that is a mixture of Fortran 77 and C and/or C++, then it requires that the macro AC_F77_LIBRARY_LDFLAGS be called in configure.in, and that either $(FLIBS) or @FLIBS@ appear in the appropriate _LDADD (for programs) or _LIBADD (for shared libraries) variables. It is the responsibility of the person writing the Makefile.am to make sure that $(FLIBS) or @FLIBS@ appears in the appropriate _LDADD or _LIBADD variable.

For example, consider the following Makefile.am:

     bin_PROGRAMS = foo
     foo_SOURCES  = main.cc foo.f
     foo_LDADD    = libfoo.la @FLIBS@
     pkglib_LTLIBRARIES = libfoo.la
     libfoo_la_SOURCES  = bar.f baz.c zardoz.cc
     libfoo_la_LIBADD   = $(FLIBS)

In this case, Automake will insist that AC_F77_LIBRARY_LDFLAGS is mentioned in configure.in. Also, if @FLIBS@ hadn’t been mentioned in foo_LDADD and libfoo_la_LIBADD, then Automake would have issued a warning.

9.10.3.1 How the Linker is Chosen

The following diagram demonstrates under what conditions a particular linker is chosen by Automake.

For example, if Fortran 77, C and C++ source code were to be compiled into a program, then the C++ linker will be used. In this case, if the C or Fortran 77 linkers required any special libraries that weren’t included by the C++ linker, then they must be manually added to an _LDADD or _LIBADD variable by the user writing the Makefile.am.

\              Linker
source      \
code        \     C        C++     Fortran
-----------------  +---------+---------+---------+
|         |         |         |
C                  |    x    |         |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C++            |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
Fortran  |         |         |    x    |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C + C++            |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C +       Fortran  |         |         |    x    |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C++ + Fortran  |         |    x    |         |
|         |         |         |
+---------+---------+---------+
|         |         |         |
C + C++ + Fortran  |         |    x    |         |
|         |         |         |
+---------+---------+---------+

9.10.4 Fortran 77 and Autoconf

The current Automake support for Fortran 77 requires a recent enough version of Autoconf that also includes support for Fortran 77. Full Fortran 77 support was added to Autoconf 2.13, so you will want to use that version of Autoconf or later.

9.11 Java Support

Automake includes support for compiled Java, using gcj, the Java front end to the GNU Compiler Collection.

Any package including Java code to be compiled must define the output variable ‘GCJ’ in configure.in; the variable ‘GCJFLAGS’ must also be defined somehow (either in configure.in or Makefile.am). The simplest way to do this is to use the AM_PROG_GCJ macro.

By default, programs including Java source files are linked with gcj.

As always, the contents of ‘AM_GCJFLAGS’ are passed to every compilation invoking gcj (in its role as an ahead-of-time compiler – when invoking it to create .class files, ‘AM_JAVACFLAGS’ is used instead). If it is necessary to pass options to gcj from Makefile.am, this variable, and not the user variable ‘GCJFLAGS’, should be used.

gcj can be used to compile .java, .class, .zip, or .jar files.

When linking, gcj requires that the main class be specified using the ‘--main=’ option. The easiest way to do this is to use the _LDFLAGS variable for the program.

9.12 Support for Other Languages

Automake currently only includes full support for C, C++ (see Section 9.8 [C++ Support], page 33), Fortran 77 (see Section 9.10 [Fortran 77 Support], page 34), and Java (see Section 9.11 [Java Support], page 38). There is only rudimentary support for other languages, support for which will be improved based on user demand.

Some limited support for adding your own languages is available via the suffix rule handling; see Section 18.2 [Suffixes], page 55.

9.13 Automatic de-ANSI-fication

Although the GNU standards allow the use of ANSI C, this can have the effect of limiting portability of a package to some older compilers (notably the SunOS C compiler).

Automake allows you to work around this problem on such machines by de-ANSI-fying each source file before the actual compilation takes place.

If the Makefile.am variable AUTOMAKE_OPTIONS (see Chapter 17 [Options], page 52) contains the option ansi2knr then code to handle de-ANSI-fication is inserted into the

generated Makefile.in.

This causes each C source file in the directory to be treated as ANSI C. If an ANSI C compiler is available, it is used. If no ANSI C compiler is available, the ansi2knr program is used to convert the source files into K&R C, which is then compiled.

The ansi2knr program is simple-minded. It assumes the source code will be formatted in a particular way; see the ansi2knr man page for details.

Support for de-ANSI-fication requires the source files ansi2knr.c and ansi2knr.1 to be in the same package as the ANSI C source; these files are distributed with Automake. Also, the package configure.in must call the macro AM_C_PROTOTYPES (see Section 5.6 [Macros], page 17).

Automake also handles finding the ansi2knr support files in some other directory in the current package. This is done by prepending the relative path to the appropriate directory to the ansi2knr option. For instance, suppose the package has ANSI C code in the src and lib subdirs. The files ansi2knr.c and ansi2knr.1 appear in lib. Then this could

appear in src/Makefile.am:
     AUTOMAKE_OPTIONS = ../lib/ansi2knr

If no directory prefix is given, the files are assumed to be in the current directory.

Files mentioned in LIBOBJS which need de-ANSI-fication will not be automatically handled. That’s because configure will generate an object name like regex.o, while make will be looking for regex_.o (when de-ANSI-fying). Eventually this problem will be fixed via autoconf magic, but for now you must put this code into your configure.in, just before

the AC_OUTPUT call:
     # This is necessary so that .o files in LIBOBJS are also built via
     # the ANSI2KNR-filtering rules.
     LIBOBJS=`echo $LIBOBJS|sed 's/\.o /\$U.o /g;s/\.o$/\$U.o/'`

Note that automatic de-ANSI-fication will not work when the package is being built for a different host architecture. That is because automake currently has no way to build ansi2knr for the build machine.

9.14 Automatic dependency tracking

As a developer it is often painful to continually update the Makefile.in whenever the include-file dependencies change in a project. Automake supplies a way to automatically track dependency changes.

Automake always uses complete dependencies for a compilation, including system headers. Automake’s model is that dependency computation should be a side effect of the build. To this end, dependencies are computed by running all compilations through a special wrapper program called depcomp. depcomp understands how to coax many different C and C++ compilers into generating dependency information in the format it requires. automake -a will install depcomp into your source tree for you. If depcomp can’t figure out how to properly invoke your compiler, dependency tracking will simply be disabled for your build.

Experience with earlier versions of Automake^{1} taught us that it is not reliable to generate dependencies only on the maintainer’s system, as configurations vary too much. So instead Automake implements dependency tracking at build time.

Automatic dependency tracking can be suppressed by putting no-dependencies in the

variable AUTOMAKE_OPTIONS, or passing no-dependencies as an argument to AM_INIT_

AUTOMAKE (this should be the prefered way). Or, you can invoke automake with the -i option. Dependency tracking is enabled by default.

The person building your package also can choose to disable dependency tracking by

configuring with --disable-dependency-tracking.

9.15 Support for executable extensions

On some platforms, such as Windows, executables are expected to have an extension such as ‘.exe’. On these platforms, some compilers (GCC among them) will automatically generate foo.exe when asked to generate foo.

1 See http://sources.redhat.com/automake/dependencies.html for more information on the history

and experiences with automatic dependency tracking in Automake

Automake provides mostly-transparent support for this. Unfortunately mostly doesn’t yet mean fully. Until the English dictionary is revised, you will have to assist Automake if your package must support those platforms.

One thing you must be aware of is that, internally, Automake rewrites something like this:

     bin_PROGRAMS = liver

to this:

     bin_PROGRAMS = liver$(EXEEXT)

The targets Automake generates are likewise given the ‘$(EXEEXT)’ extension. EXEEXT However, Automake cannot apply this rewriting to configure substitutions. This means

that if you are conditionally building a program using such a substitution, then your configure.in must take care to add ‘$(EXEEXT)’ when constructing the output variable.

With Autoconf 2.13 and earlier, you must explicitly use AC_EXEEXT to get this support. With Autoconf 2.50, AC_EXEEXT is run automatically if you configure a compiler (say,

through AC_PROG_CC).

Sometimes maintainers like to write an explicit link rule for their program. Without executable extension support, this is easy—you simply write a target with the same name as the program. However, when executable extension support is enabled, you must instead add the ‘$(EXEEXT)’ suffix.

Unfortunately, due to the change in Autoconf 2.50, this means you must always add this extension. However, this is a problem for maintainers who know their package will never run on a platform that has executable extensions. For those maintainers, the no-exeext option (see Chapter 17 [Options], page 52) will disable this feature. This works in a fairly ugly way; if no-exeext is seen, then the presence of a target named foo in Makefile.am will override an automake-generated target of the form foo$(EXEEXT). Without the no-exeext option, this use will give an error.

10 Other Derived Objects

Automake can handle derived objects which are not C programs. Sometimes the support for actually building such objects must be explicitly supplied, but Automake will still automatically handle installation and distribution.

10.1 Executable Scripts

It is possible to define and install programs which are scripts. Such programs are listed using the ‘SCRIPTS’ primary name. Automake doesn’t define any dependencies for scripts; the Makefile.am should include the appropriate rules.

Automake does not assume that scripts are derived objects; such objects must be deleted by hand (see Chapter 14 [Clean], page 48).

The automake program itself is a Perl script that is generated at configure time from automake.in. Here is how this is handled:

     bin_SCRIPTS = automake

Since automake appears in the AC_OUTPUT macro, a target for it is automatically generated, and it is also automatically cleaned (despite the fact it’s a script).

Script objects can be installed in bindir, sbindir, libexecdir, or pkgdatadir.

Scripts that need not being installed can be listed in noinst_SCRIPTS, and among them, those which are needed only by make check should go in check_SCRIPTS.

10.2 Header files

Header files are specified by the ‘HEADERS’ family of variables. Generally header files are not installed, so the noinst_HEADERS variable will be the most used.^{2}

All header files must be listed somewhere; missing ones will not appear in the distribution. Often it is clearest to list uninstalled headers with the rest of the sources for a program. See Section 9.1 [A Program], page 23. Headers listed in a ‘_SOURCES’ variable need not be listed in any ‘_HEADERS’ variable.

   Headers can be installed in includedir, oldincludedir, or pkgincludedir.

10.3 Architecture-independent data files

Automake supports the installation of miscellaneous data files using the ‘DATA’ family of variables.

Such data can be installed in the directories datadir, sysconfdir, sharedstatedir,

localstatedir, or pkgdatadir.

By default, data files are not included in a distribution. Of course, you can use the ‘dist_’ prefix to change this on a per-variable basis.

Here is how Automake declares its auxiliary data files:

     dist_pkgdata_DATA = clean-kr.am clean.am ...

10.4 Built sources

Occasionally a file which would otherwise be called ‘source’ (e.g. a C ‘.h’ file) is actually derived from some other file. Such files should be listed in the BUILT_SOURCES variable.

BUILT_SOURCES is actually a bit of a misnomer, as any file which must be created early in the build process can be listed in this variable.

A source file listed in BUILT_SOURCES is created before the other all targets are made. However, such a source file is not compiled unless explicitly requested by mentioning it in some other ‘_SOURCES’ variable.

So, for instance, if you had header files which were created by a script run at build time, then you would list these headers in BUILT_SOURCES, to ensure that they would be built before any other compilations (perhaps ones using these headers) were started.

2 However, for the case of a non-installed header file that is actually used by a particular program, we

recommend listing it in the program’s ‘_SOURCES’ variable instead of in noinst_HEADERS. We believe this

is more clear.

11 Other GNU Tools

Since Automake is primarily intended to generate Makefile.ins for use in GNU programs, it tries hard to interoperate with other GNU tools.

11.1 Emacs Lisp

Automake provides some support for Emacs Lisp. The ‘LISP’ primary is used to hold a list of .el files. Possible prefixes for this primary are ‘lisp_’ and ‘noinst_’. Note that

if lisp_LISP is defined, then configure.in must run AM_PATH_LISPDIR (see Section 5.6

[Macros], page 17).

By default Automake will byte-compile all Emacs Lisp source files using the Emacs found by AM_PATH_LISPDIR. If you wish to avoid byte-compiling, simply define the variable ELCFILES to be empty. Byte-compiled Emacs Lisp files are not portable among all versions of Emacs, so it makes sense to turn this off if you expect sites to have more than one version of Emacs installed. Furthermore, many packages don’t actually benefit from bytecompilation. Still, we recommend that you leave it enabled by default. It is probably better for sites with strange setups to cope for themselves than to make the installation less nice for everybody else.

11.2 Gettext

If AM_GNU_GETTEXT is seen in configure.in, then Automake turns on support for GNU gettext, a message catalog system for internationalization (see Section “GNU Gettext” in GNU gettext utilities).

The gettext support in Automake requires the addition of two subdirectories to the package, intl and po. Automake insures that these directories exist and are mentioned in

SUBDIRS.

11.3 Libtool

Automake provides support for GNU Libtool (see Section “Introduction” in The Libtool Manual) with the ‘LTLIBRARIES’ primary. See Section 9.3 [A Shared Library], page 27.

11.4 Java

Automake provides some minimal support for Java compilation with the ‘JAVA’ primary.

Any .java files listed in a ‘_JAVA’ variable will be compiled with JAVAC at build time. By default, .class files are not included in the distribution.

Currently Automake enforces the restriction that only one ‘_JAVA’ primary can be used in a given Makefile.am. The reason for this restriction is that, in general, it isn’t possible to know which .class files were generated from which .java files – so it would be impossible to know which files to install where. For instance, a .java file can define multiple classes; the resulting .class file names cannot be predicted without parsing the .java file.

There are a few variables which are used when compiling Java sources:

JAVAC The name of the Java compiler. This defaults to ‘javac’.

JAVACFLAGS

The flags to pass to the compiler. This is considered to be a user variable (see Section 2.5 [User Variables], page 4).

AM_JAVACFLAGS

More flags to pass to the Java compiler. This, and not JAVACFLAGS, should be used when it is necessary to put Java compiler flags into Makefile.am.

JAVAROOT The value of this variable is passed to the ‘-d’ option to javac. It defaults to

          ‘$(top_builddir)’.
CLASSPATH_ENV

This variable is an sh expression which is used to set the CLASSPATH environment variable on the javac command line. (In the future we will probably handle class path setting differently.)

11.5 Python

Automake provides support for Python compilation with the ‘PYTHON’ primary.

Any files listed in a ‘_PYTHON’ variable will be byte-compiled with py-compile at install time. py-compile actually creates both standard (.pyc) and byte-compiled (.pyo) versions of the source files. Note that because byte-compilation occurs at install time, any files listed in ‘noinst_PYTHON’ will not be compiled. Python source files are included in the distribution by default.

Automake ships with an Autoconf macro called AM_PATH_PYTHON which will determine some Python-related directory variables (see below). If you have called AM_PATH_ PYTHON from configure.in, then you may use the following variables to list you Python

source files in your variables: ‘python_PYTHON’, ‘pkgpython_PYTHON’, ‘pyexecdir_PYTHON’,

‘pkgpyexecdir_PYTHON’, depending where you want your files installed.

AM_PATH_PYTHON takes a single optional argument. This argument, if present, is the minimum version of Python which can be used for this package. If the version of Python found on the system is older than the required version, then AM_PATH_PYTHON will cause an error.

AM_PATH_PYTHON creates several output variables based on the Python installation found during configuration.

PYTHON The name of the Python executable.

PYTHON_VERSION

The Python version number, in the form major.minor (e.g. ‘1.5’). This is currently the value of sys.version[:3].

PYTHON_PREFIX

The string $prefix. This term may be used in future work which needs the contents of Python’s sys.prefix, but general consensus is to always use the value from configure.

PYTHON_EXEC_PREFIX

The string $exec_prefix. This term may be used in future work which needs the contents of Python’s sys.exec_prefix, but general consensus is to always use the value from configure.

PYTHON_PLATFORM

The canonical name used by Python to describe the operating system, as given by sys.platform. This value is sometimes needed when building Python extensions.

pythondir

The directory name for the site-packages subdirectory of the standard Python install tree.

pkgpythondir

This is is the directory under pythondir which is named after the package. That is, it is ‘$(pythondir)/$(PACKAGE)’. It is provided as a convenience.

pyexecdir

This is the directory where Python extension modules (shared libraries) should be installed.

pkgpyexecdir

This is a convenience variable which is defined as ‘$(pyexecdir)/$(PACKAGE)’.

12 Building documentation

Currently Automake provides support for Texinfo and man pages.

12.1 Texinfo

If the current directory contains Texinfo source, you must declare it with the ‘TEXINFOS’ primary. Generally Texinfo files are converted into info, and thus the info_TEXINFOS variable is most commonly used here. Any Texinfo source file must end in the .texi, .txi, or .texinfo extension. We recommend .texi for new manuals.

Automake generates rules to build .info, .dvi, .ps, and .pdf files from your Texinfo sources. The .info files are built by make all and installed by make install (unless you use no-installinfo, see below). The other files can be built on request by make dvi, make

ps, and make pdf.

If the .texi file @includes version.texi, then that file will be automatically generated. The file version.texi defines four Texinfo flag you can reference using @value{EDITION},

@value{VERSION}, @value{UPDATED}, and @value{UPDATED-MONTH}.
EDITION

VERSION Both of these flags hold the version number of your program. They are kept separate for clarity.

UPDATED This holds the date the primary .texi file was last modified.

UPDATED-MONTH

This holds the name of the month in which the primary .texi file was last modified.

The version.texi support requires the mdate-sh program; this program is supplied with Automake and automatically included when automake is invoked with the --add-

missing option.

If you have multiple Texinfo files, and you want to use the version.texi feature, then you have to have a separate version file for each Texinfo file. Automake will treat any include in a Texinfo file that matches ‘vers*.texi’ just as an automatically generated version file.

When an info file is rebuilt, the program named by the MAKEINFO variable is used to invoke it. If the makeinfo program is found on the system then it will be used by default; otherwise missing will be used instead. The flags in the variables MAKEINFOFLAGS and AM_ MAKEINFOFLAGS will be passed to the makeinfo invocation; the first of these is intended for use by the user (see Section 2.5 [User Variables], page 4) and the second by the Makefile.am writer.

Sometimes an info file actually depends on more than one .texi file. For instance, in GNU Hello, hello.texi includes the file gpl.texi. You can tell Automake about these dependencies using the texi_TEXINFOS variable. Here is how GNU Hello does it:

     info_TEXINFOS = hello.texi
     hello_TEXINFOS = gpl.texi

By default, Automake requires the file texinfo.tex to appear in the same directory as the Texinfo source. However, if you used AC_CONFIG_AUX_DIR in configure.in (see Section “Finding ‘configure’ Input” in The Autoconf Manual), then texinfo.tex is looked for there. Automake supplies texinfo.tex if ‘--add-missing’ is given.

If your package has Texinfo files in many directories, you can use the variable TEXINFO_ TEX to tell Automake where to find the canonical texinfo.tex for your package. The value of this variable should be the relative path from the current Makefile.am to texinfo.tex:

     TEXINFO_TEX = ../doc/texinfo.tex

The option ‘no-texinfo.tex’ can be used to eliminate the requirement for texinfo.tex. Use of the variable TEXINFO_TEX is preferable, however, because that allows the dvi, ps, and pdf targets to still work.

Automake generates an install-info target; some people apparently use this. By default, info pages are installed by ‘make install’. This can be prevented via the

no-installinfo option.

12.2 Man pages

A package can also include man pages (but see the GNU standards on this matter, Section “Man Pages” in The GNU Coding Standards.) Man pages are declared using the ‘MANS’ primary. Generally the man_MANS variable is used. Man pages are automatically installed in the correct subdirectory of mandir, based on the file extension.

File extensions such as ‘.1c’ are handled by looking for the valid part of the extension and using that to determine the correct subdirectory of mandir. Valid section names are the digits ‘0’ through ‘9’, and the letters ‘l’ and ‘n’.

Sometimes developers prefer to name a man page something like foo.man in the source, and then rename it to have the correct suffix, e.g. foo.1, when installing the file. Automake also supports this mode. For a valid section named SECTION, there is a corresponding directory named ‘manSECTIONdir’, and a corresponding ‘_MANS’ variable. Files listed in such a variable are installed in the indicated section. If the file already has a valid suffix, then it is installed as-is; otherwise the file suffix is changed to match the section.

For instance, consider this example:

     man1_MANS = rename.man thesame.1 alsothesame.1c

In this case, rename.man will be renamed to rename.1 when installed, but the other files will keep their names.

By default, man pages are installed by ‘make install’. However, since the GNU project does not require man pages, many maintainers do not expend effort to keep the man pages up to date. In these cases, the no-installman option will prevent the man pages from being installed by default. The user can still explicitly install them via ‘make install-man’.

Here is how the man pages are handled in GNU cpio (which includes both Texinfo documentation and man pages):

     man_MANS = cpio.1 mt.1
     EXTRA_DIST = $(man_MANS)

Man pages are not currently considered to be source, because it is not uncommon for man pages to be automatically generated. Therefore they are not automatically included in the distribution. However, this can be changed by use of the ‘dist_’ prefix.

The ‘nobase_’ prefix is meaningless for man pages and is disallowed.

13 What Gets Installed

13.1 Basics of installation

Naturally, Automake handles the details of actually installing your program once it has been built. All files named by the various primaries are automatically installed in the appropriate places when the user runs make install.

A file named in a primary is installed by copying the built file into the appropriate directory. The base name of the file is used when installing.

     bin_PROGRAMS = hello subdir/goodbye

In this example, both ‘hello’ and ‘goodbye’ will be installed in $(bindir).

Sometimes it is useful to avoid the basename step at install time. For instance, you might have a number of header files in subdirectories of the source tree which are laid out precisely how you want to install them. In this situation you can use the ‘nobase_’ prefix to suppress the base name step. For example:

     nobase_include_HEADERS = stdio.h sys/types.h
   Will install stdio.h in $(includedir) and types.h in $(includedir)/sys.

13.2 The two parts of install

Automake generates separate install-data and install-exec targets, in case the installer is installing on multiple machines which share directory structure—these targets allow the machine-independent parts to be installed only once. install-exec installs platformdependent files, and install-data installs platform-independent files. The install target depends on both of these targets. While Automake tries to automatically segregate objects into the correct category, the Makefile.am author is, in the end, responsible for making sure this is done correctly.

Variables using the standard directory prefixes ‘data’, ‘info’, ‘man’, ‘include’,

‘oldinclude’, ‘pkgdata’, or ‘pkginclude’ (e.g.  ‘data_DATA’) are installed by
‘install-data’.

Variables using the standard directory prefixes ‘bin’, ‘sbin’, ‘libexec’, ‘sysconf’,

‘localstate’, ‘lib’, or ‘pkglib’ (e.g. ‘bin_PROGRAMS’) are installed by ‘install-exec’.

Any variable using a user-defined directory prefix with ‘exec’ in the name (e.g. ‘myexecbin_PROGRAMS’ is installed by ‘install-exec’. All other user-defined prefixes are installed by ‘install-data’.

13.3 Extending installation

It is possible to extend this mechanism by defining an install-exec-local or install-data-local target. If these targets exist, they will be run at ‘make install’ time. These rules can do almost anything; care is required.

Automake also supports two install hooks, install-exec-hook and install-data-hook. These hooks are run after all other install rules of the appropriate type, exec or data, have completed. So, for instance, it is possible to perform post-installation modifications using an install hook.

13.4 Staged installs

Automake generates support for the ‘DESTDIR’ variable in all install rules. ‘DESTDIR’ is used during the ‘make install’ step to relocate install objects into a staging area. Each object and path is prefixed with the value of ‘DESTDIR’ before being copied into the install area. Here is an example of typical DESTDIR usage:

     make DESTDIR=/tmp/staging install

This places install objects in a directory tree built under /tmp/staging. If /gnu/bin/foo and /gnu/share/aclocal/foo.m4 are to be installed, the above command would install

/tmp/staging/gnu/bin/foo and /tmp/staging/gnu/share/aclocal/foo.m4.

This feature is commonly used to build install images and packages. For more information, see Section “Makefile Conventions” in The GNU Coding Standards.

Support for ‘DESTDIR’ is implemented by coding it directly into the install rules. If your Makefile.am uses a local install rule (e.g., install-exec-local) or an install hook, then you must write that code to respect ‘DESTDIR’.

13.5 Rules for the user

Automake also generates an uninstall target, an installdirs target, and an

install-strip target.

Automake supports uninstall-local and uninstall-hook. There is no notion of separate uninstalls for “exec” and “data”, as these features would not provide additional functionality.

Note that uninstall is not meant as a replacement for a real packaging tool.

14 What Gets Cleaned

The GNU Makefile Standards specify a number of different clean rules. See See Section “Standard Targets for Users” in The GNU Coding Standards.

Generally the files that can be cleaned are determined automatically by Automake. Of course, Automake also recognizes some variables that can be defined to specify additional

files to clean. These variables are MOSTLYCLEANFILES, CLEANFILES, DISTCLEANFILES, and
MAINTAINERCLEANFILES.

As the GNU Standards aren’t always explicit as to which files should be removed by which target, we’ve adopted a heuristic which we believe was first formulated by François Pinard:

• If make built it, and it is commonly something that one would want to rebuild (for instance, a .o file), then mostlyclean should delete it.

• Otherwise, if make built it, then clean should delete it.

• If configure built it, then distclean should delete it

• If the maintainer built it, then maintainer-clean should delete it.

We recommend that you follow this same set of heuristics in your Makefile.am.

15 What Goes in a Distribution

15.1 Basics of distribution

The dist target in the generated Makefile.in can be used to generate a gzip’d tar file and other flavors of archive for distribution. The files is named based on the ‘PACKAGE’ and ‘VERSION’ variables defined by AM_INIT_AUTOMAKE (see Section 5.6 [Macros], page 17); more precisely the gzip’d tar file is named ‘package-version.tar.gz’. You can use the make variable ‘GZIP_ENV’ to control how gzip is run. The default setting is ‘--best’.

For the most part, the files to distribute are automatically found by Automake: all source files are automatically included in a distribution, as are all Makefile.ams and Makefile.ins. Automake also has a built-in list of commonly used files which are automatically included if they are found in the current directory (either physically, or as the target of a Makefile.am rule). This list is printed by ‘automake --help’. Also, files which are read by configure (i.e. the source files corresponding to the files specified in various Autoconf macros such as AC_CONFIG_FILES and siblings) are automatically distributed.

Still, sometimes there are files which must be distributed, but which are not covered in the automatic rules. These files should be listed in the EXTRA_DIST variable. You can mention files from subdirectories in EXTRA_DIST.

You can also mention a directory in EXTRA_DIST; in this case the entire directory will be recursively copied into the distribution. Please note that this will also copy everything in the directory, including CVS/RCS version control files. We recommend against using this feature.

15.2 Fine-grained distribution control

Sometimes you need tighter control over what does not go into the distribution; for instance you might have source files which are generated and which you do not want to distribute. In this case Automake gives fine-grained control using the ‘dist’ and ‘nodist’ prefixes. Any primary or ‘_SOURCES’ variable can be prefixed with ‘dist_’ to add the listed files to the distribution. Similarly, ‘nodist_’ can be used to omit the files from the distribution.

As an example, here is how you would cause some data to be distributed while leaving some source code out of the distribution:

     dist_data_DATA = distribute-this
     bin_PROGRAMS = foo
     nodist_foo_SOURCES = do-not-distribute.c

15.3 The dist hook

Another way to to use this is for removing unnecessary files that get recursively included by specifying a directory in EXTRA DIST:

     EXTRA_DIST = doc
     dist-hook:
     rm -rf `find $(distdir)/doc -name CVS`

If you define SUBDIRS, Automake will recursively include the subdirectories in the distribution. If SUBDIRS is defined conditionally (see Chapter 20 [Conditionals], page 56), Automake will normally include all directories that could possibly appear in SUBDIRS in the distribution. If you need to specify the set of directories conditionally, you can set the variable DIST_SUBDIRS to the exact list of subdirectories to include in the distribution.

Occasionally it is useful to be able to change the distribution before it is packaged up. If the dist-hook target exists, it is run after the distribution directory is filled, but before the actual tar (or shar) file is created. One way to use this is for distributing files in subdirectories for which a new Makefile.am is overkill:

     dist-hook:
             mkdir $(distdir)/random
             cp -p $(srcdir)/random/a1 $(srcdir)/random/a2 $(distdir)/random

15.4 Checking the distribution

Automake also generates a distcheck target which can be of help to ensure that a given distribution will actually work. distcheck makes a distribution, then tries to do a VPATH build, run the testsuite, and finally make another tarfile to ensure the distribution is selfcontained.

Building the package involves running ./configure. If you need to supply additional flags to configure, define them in the DISTCHECK_CONFIGURE_FLAGS variable, either in your top-level Makefile.am, or on the commande line when invoking make.

If the target distcheck-hook is defined in your Makefile.am, then it will be invoked by distcheck after the new distribution has been unpacked, but before the unpacked copy is configured and built. Your distcheck-hook can do almost anything, though as always caution is advised. Generally this hook is used to check for potential distribution errors not caught by the standard mechanism.

Speaking about potential distribution errors, distcheck will also ensure that the distclean target actually removes all built files. This is done by running make

distcleancheck at the end of the VPATH build. By default, distcleancheck will

run distclean and then make sure the build tree has been emptied by running $(distcleancheck_listfiles). Usually this check will find generated files that you forgot to add to the DISTCLEANFILES variable (see Chapter 14 [Clean], page 48).

The distcleancheck behaviour should be ok for most packages, otherwise you have the possibility to override the definitition of either the distcleancheck target, or the

$(distcleancheck_listfiles) variable. For instance to disable distcleancheck com-

pletely, add the following rule to your top-level Makefile.am:

     distcleancheck:
             @:

If you want distcleancheck to ignore built files which have not been cleaned because they are also part of the distribution, add the following definition instead:

     distcleancheck_listfiles = \
       find -type f -exec sh -c 'test -f $(scrdir)/{} || echo {}'

The above definition is not the default because it’s usually an error if your Makefiles cause some distributed files to be rebuilt when the user build the package. (Think about the user missing the tool required to build the file; or if the required tool is built by your package, consider the cross-compilation case where it can’t be run.)

distcheck also checks that the uninstall target works properly, both for ordinary and ‘DESTDIR’ builds. It does this by invoking make uninstall, and then it checks the install tree to see if any files are left over. This check will make sure that you correctly coded your uninstall-related targets.

By default, the checking is done by the distuninstallcheck target, and the list of files in the install tree is generated by $(distuninstallcheck_listfiles) (this is a variable whose value is a shell command to run that prints the list of files to stdout).

Either of these can be overridden to modify the behavior of distcheck. For instance, to disable this check completely, you would write:

     distuninstallcheck:
             @:

15.5 The types of distributions

Automake generates a ‘.tar.gz’ file when asked to create a distribution and other archives formats, Chapter 17 [Options], page 52. The target dist-gzip generates the ‘.tar.gz’ file only.

16 Support for test suites

Automake supports two forms of test suites.

16.1 Simple Tests

If the variable TESTS is defined, its value is taken to be a list of programs to run in order to do the testing. The programs can either be derived objects or source objects; the generated rule will look both in srcdir and .. Programs needing data files should look for them in srcdir (which is both an environment variable and a make variable) so they work when building in a separate directory (see Section “Build Directories ” in The Autoconf Manual), and in particular for the distcheck target (see Chapter 15 [Dist], page 48).

The number of failures will be printed at the end of the run. If a given test program exits with a status of 77, then its result is ignored in the final count. This feature allows non-portable tests to be ignored in environments where they don’t make sense.

The variable TESTS_ENVIRONMENT can be used to set environment variables for the test run; the environment variable srcdir is set in the rule. If all your test programs are scripts, you can also set TESTS_ENVIRONMENT to an invocation of the shell (e.g. ‘$(SHELL) -x’); this can be useful for debugging the tests.

You may define the variable XFAIL_TESTS to a list of tests (usually a subset of TESTS) that are expected to fail. This will reverse the result of those tests.

Automake ensures that each program listed in TESTS is built before any tests are run; you can list both source and derived programs in TESTS. For instance, you might want to run a C program as a test. To do this you would list its name in TESTS and also in check_PROGRAMS, and then specify it as you would any other program.

16.2 DejaGNU Tests

If ‘dejagnu’ (ftp://ftp.gnu.org/gnu/dejagnu/) appears in AUTOMAKE_OPTIONS, then a

dejagnu-based test suite is assumed. The variable DEJATOOL is a list of names which are passed, one at a time, as the --tool argument to runtest invocations; it defaults to the name of the package.

The variable RUNTESTDEFAULTFLAGS holds the --tool and --srcdir flags that are passed to dejagnu by default; this can be overridden if necessary.

The variables EXPECT and RUNTEST can also be overridden to provide project-specific values. For instance, you will need to do this if you are testing a compiler toolchain, because the default values do not take into account host and target names.

The contents of the variable RUNTESTFLAGS are passed to the runtest invocation. This is considered a “user variable” (see Section 2.5 [User Variables], page 4). If you need to set

runtest flags in Makefile.am, you can use AM_RUNTESTFLAGS instead.

Automake will generate rules to create a local site.exp file, defining various variables detected by ./configure. This file is automatically read by DejaGnu. It is ok for the user of a package to edit this file in order to tune the test suite. However this is not the place where the test suite author should define new variables: this should be done elsewhere in the real test suite code. Especially, site.exp should not be distributed.

In either case, the testing is done via ‘make check’.

16.3 Install Tests

The installcheck target is available to the user as a way to run any tests after the package has been installed. You can add tests to this by writing an installcheck-local target.

17 Changing Automake’s Behavior

Various features of Automake can be controlled by options in the Makefile.am. Such options are applied on a per-Makefile basis when listed in a special Makefile variable named AUTOMAKE_OPTIONS. They are applied globally to all processed Makefiles when listed in the first argument of AM_INIT_AUTOMAKE in configure.in. Currently understood options are:

gnits
gnu
foreign
cygnus

Set the strictness as appropriate. The gnits option also implies readme-alpha

and check-news.
ansi2knr
path/ansi2knr

Turn on automatic de-ANSI-fication. See Section 9.13 [ANSI], page 38. If preceded by a path, the generated Makefile.in will look in the specified directory to find the ansi2knr program. The path should be a relative path to another directory in the same distribution (Automake currently does not check this).

check-news

Cause make dist to fail unless the current version number appears in the first few lines of the NEWS file.

dejagnu Cause dejagnu-specific rules to be generated. See Chapter 16 [Tests], page 50.

dist-bzip2

Generate a dist-bzip2 target, creating a bzip2 tar archive of the distribution. dist will create it in addition to the other formats. bzip2 archives are frequently smaller than gzipped archives.

dist-shar

Generate a dist-shar target, creating a shar archive of the distribution. dist will create it in addition to the other formats.

dist-zip Generate a dist-zip target, creating a zip archive of the distribution. dist will create it in addition to the other formats.

dist-tarZ

Generate a dist-tarZ target, creating a compressed tar archive of the distribution. dist will create it in addition to the other formats.

no-define

This options is meaningful only when passed as an argument to AM_INIT_ AUTOMAKE. It will prevent the PACKAGE and VERSION variables to be AC_DEFINEd.

no-dependencies

This is similar to using ‘--include-deps’ on the command line, but is useful for those situations where you don’t have the necessary bits to make automatic dependency tracking work See Section 9.14 [Dependencies], page 39. In this case the effect is to effectively disable automatic dependency tracking.

no-exeext

If your Makefile.am defines a target ‘foo’, it will override a target named ‘foo$(EXEEXT)’. This is necessary when EXEEXT is found to be empty. However, by default automake will generate an error for this use. The no-exeext option will disable this error. This is intended for use only where it is known in advance that the package will not be ported to Windows, or any other operating system using extensions on executables.

no-installinfo

The generated Makefile.in will not cause info pages to be built or installed by default. However, info and install-info targets will still be available. This option is disallowed at ‘GNU’ strictness and above.

no-installman

The generated Makefile.in will not cause man pages to be installed by default. However, an install-man target will still be available for optional installation. This option is disallowed at ‘GNU’ strictness and above.

nostdinc This option can be used to disable the standard ‘-I’ options which are ordinarily automatically provided by Automake.

no-texinfo.tex

Don’t require texinfo.tex, even if there are texinfo files in this directory.

readme-alpha

If this release is an alpha release, and the file README-alpha exists, then it will be added to the distribution. If this option is given, version numbers are expected to follow one of two forms. The first form is ‘MAJOR.MINOR.ALPHA’, where each element is a number; the final period and number should be left off for non-alpha releases. The second form is ‘MAJOR.MINORALPHA’, where ALPHA is a letter; it should be omitted for non-alpha releases.

std-options

Make the installcheck target check that installed scripts and programs support the --help and --version options. This also provides a basic check that the program’s run-time dependencies are satisfied after installation. In a few situations, programs (or scripts) have to be exempted from this test. For instance false (from GNU sh-utils) is never successful, even for --help or --version. You can list such programs in the variable AM_INSTALLCHECK_

STD_OPTIONS_EXEMPT.
subdir-objects

If this option is specified, then objects are placed into the subdirectory of the build directory corresponding to the subdirectory of the source file. For instance if the source file is subdir/file.cxx, then the output file would be

subdir/file.o.

version A version number (e.g. ‘0.30’) can be specified. If Automake is not newer than the version specified, creation of the Makefile.in will be suppressed.

-Wcategory or --warnings=category

These options behave exactly like their command-line counterpart (see Chapter 4 [Invoking Automake], page 9). This allows you to enable or disable some warning categories on a per-file basis. You can also setup some warnings for your entire project; for instance try AM_INIT_AUTOMAKE([-Wall]) in

          your configure.in.

Unrecognized options are diagnosed by automake.

If you want an option to apply to all the files in the tree, you can use the AM_INIT_ AUTOMAKE macro in configure.in. See Section 5.6 [Macros], page 17.

18 Miscellaneous Rules

There are a few rules and variables that didn’t fit anywhere else.

18.1 Interfacing to etags

Automake will generate rules to generate TAGS files for use with GNU Emacs under some circumstances.

If any C, C++ or Fortran 77 source code or headers are present, then tags and TAGS targets will be generated for the directory.

At the topmost directory of a multi-directory package, a tags target file will be generated which, when run, will generate a TAGS file that includes by reference all TAGS files from subdirectories.

The tags target will also be generated if the variable ETAGS_ARGS is defined. This variable is intended for use in directories which contain taggable source that etags does not understand. The user can use the ETAGSFLAGS to pass additional flags to etags; AM_ ETAGSFLAGS is also available for use in Makefile.am.

Here is how Automake generates tags for its source, and for nodes in its Texinfo file:

     ETAGS_ARGS = automake.in --lang=none \
      --regex='/^@node[ \t]+\([^,]+\)/\1/' automake.texi

If you add filenames to ‘ETAGS_ARGS’, you will probably also want to set ‘TAGS_DEPENDENCIES’. The contents of this variable are added directly to the dependencies for the tags target.

Automake also generates a ctags target which can be used to build vi-style tags files. The variable CTAGS is the name of the program to invoke (by default ‘ctags’); CTAGSFLAGS can be used by the user to pass additional flags, and AM_CTAGSFLAGS can be used by the

Makefile.am.

Automake will also generate an ID target which will run mkid on the source. This is only supported on a directory-by-directory basis.

Automake also supports the GNU Global Tags program (http://www.gnu.org/ software/global/). The GTAGS target runs Global Tags automatically and puts the result in the top build directory. The variable GTAGS_ARGS holds arguments which are passed to

gtags.

18.2 Handling new file extensions

It is sometimes useful to introduce a new implicit rule to handle a file type that Automake does not know about.

For instance, suppose you had a compiler which could compile ‘.foo’ files to ‘.o’ files. You would simply define an suffix rule for your language:

     .foo.o:
             foocc -c -o $@ $<

Then you could directly use a ‘.foo’ file in a ‘_SOURCES’ variable and expect the correct results:

     bin_PROGRAMS = doit
     doit_SOURCES = doit.foo

This was the simpler and more common case. In other cases, you will have to help Automake to figure which extensions you are defining your suffix rule for. This usually happens when your extensions does not start with a dot. Then, all you have to do is to put a list of new suffixes in the SUFFIXES variable before you define your implicit rule.

For instance the following definition prevents Automake to misinterpret ‘.idlC.cpp:’ as an attemp to transform ‘.idlC’ into ‘.cpp’.

     SUFFIXES = .idl C.cpp
     .idlC.cpp:
             # whatever

As you may have noted, the SUFFIXES variable behaves like the .SUFFIXES special target of make. You should not touch .SUFFIXES yourself, but use SUFFIXES instead and let Automake generate the suffix list for .SUFFIXES. Any given SUFFIXES go at the start of the generated suffixes list, followed by Automake generated suffixes not already in the list.

18.3 Support for Multilibs

Automake has support for an obscure feature called multilibs. A multilib is a library which is built for multiple different ABIs at a single time; each time the library is built with a different target flag combination. This is only useful when the library is intended to be cross-compiled, and it is almost exclusively used for compiler support libraries.

The multilib support is still experimental. Only use it if you are familiar with multilibs and can debug problems you might encounter.

19 Include

Automake supports an include directive which can be used to include other Makefile fragments when automake is run. Note that these fragments are read and interpreted by automake, not by make. As with conditionals, make has no idea that include is in use.

There are two forms of include:

include $(srcdir)/file

Include a fragment which is found relative to the current source directory.

include $(top_srcdir)/file

Include a fragment which is found relative to the top source directory.

Note that if a fragment is included inside a conditional, then the condition applies to the entire contents of that fragment.

Makefile fragments included this way are always distributed because there are needed to

rebuild Makefile.in.

20 Conditionals

Automake supports a simple type of conditionals.

Before using a conditional, you must define it by using AM_CONDITIONAL in the configure.in file (see Section 5.6 [Macros], page 17).

[Macro] AM_CONDITIONAL (conditional, condition)

The conditional name, conditional, should be a simple string starting with a letter and containing only letters, digits, and underscores. It must be different from ‘TRUE’ and ‘FALSE’ which are reserved by Automake. The shell condition (suitable for use in a shell if statement) is evaluated when configure is run. Note that you must arrange for every AM_CONDITIONAL to be invoked every time configure is run – if AM_CONDITIONAL is run conditionally (e.g., in a shell if statement), then the result will confuse automake.

Conditionals typically depend upon options which the user provides to the configure script. Here is an example of how to write a conditional which is true if the user uses the

‘--enable-debug’ option.
AC_ARG_ENABLE(debug,
[  --enable-debug    Turn on debugging],
[case "${enableval}" in
  yes) debug=true ;;
  no)  debug=false ;;
  *) AC_MSG_ERROR(bad value ${enableval} for --enable-debug) ;;
esac],[debug=false])
AM_CONDITIONAL(DEBUG, test x$debug = xtrue)

Here is an example of how to use that conditional in Makefile.am:

if DEBUG
DBG = debug
else
DBG =
endif
noinst_PROGRAMS = $(DBG)

This trivial example could also be handled using EXTRA PROGRAMS (see Section 9.1.4 [Conditional Programs], page 26).

You may only test a single variable in an if statement, possibly negated using ‘!’. The else statement may be omitted. Conditionals may be nested to any depth. You may specify an argument to else in which case it must be the negation of the condition used for the current if. Similarly you may specify the condition which is closed by an end:

     if DEBUG
     DBG = debug
     else !DEBUG
     DBG =
     endif !DEBUG

Unbalanced conditions are errors.

Note that conditionals in Automake are not the same as conditionals in GNU Make. Automake conditionals are checked at configure time by the configure script, and affect the translation from Makefile.in to Makefile. They are based on options passed to configure and on results that configure has discovered about the host system. GNU Make conditionals are checked at make time, and are based on variables passed to the make program or defined in the Makefile.

Automake conditionals will work with any make program.

21 The effect of --gnu and --gnits

The ‘--gnu’ option (or ‘gnu’ in the ‘AUTOMAKE_OPTIONS’ variable) causes automake to check the following:

  • The files INSTALL, NEWS, README, AUTHORS, and ChangeLog, plus one of COPYING.LIB,

COPYING.LESSER or COPYING, are required at the topmost directory of the package.

• The options ‘no-installman’ and ‘no-installinfo’ are prohibited.

Note that this option will be extended in the future to do even more checking; it is advisable to be familiar with the precise requirements of the GNU standards. Also, ‘--gnu’ can require certain non-standard GNU programs to exist for use by various maintainer-only targets; for instance in the future pathchk might be required for ‘make dist’.

The ‘--gnits’ option does everything that ‘--gnu’ does, and checks the following as well:

• ‘make installcheck’ will check to make sure that the --help and --version really print a usage message and a version string, respectively. This is the std-options option (see Chapter 17 [Options], page 52).

• ‘make dist’ will check to make sure the NEWS file has been updated to the current version.

• ‘VERSION’ is checked to make sure its format complies with Gnits standards.

• If ‘VERSION’ indicates that this is an alpha release, and the file README-alpha appears in the topmost directory of a package, then it is included in the distribution. This is done in ‘--gnits’ mode, and no other, because this mode is the only one where version number formats are constrained, and hence the only mode where Automake can automatically determine whether README-alpha should be included.

• The file THANKS is required.

22 The effect of --cygnus

Some packages, notably GNU GCC and GNU gdb, have a build environment originally written at Cygnus Support (subsequently renamed Cygnus Solutions, and then later purchased by Red Hat). Packages with this ancestry are sometimes referred to as “Cygnus” trees.

A Cygnus tree has slightly different rules for how a Makefile.in is to be constructed. Passing ‘--cygnus’ to automake will cause any generated Makefile.in to comply with Cygnus rules.

Here are the precise effects of ‘--cygnus’:

• Info files are always created in the build directory, and not in the source directory.

• texinfo.tex is not required if a Texinfo source file is specified. The assumption is that the file will be supplied, but in a place that Automake cannot find. This assumption is an artifact of how Cygnus packages are typically bundled.

• ‘make dist’ is not supported, and the rules for it are not generated. Cygnus-style trees use their own distribution mechanism.

• Certain tools will be searched for in the build tree as well as in the user’s ‘PATH’. These

    tools are runtest, expect, makeinfo and texi2dvi.

• --foreign is implied.

  • The options ‘no-installinfo’ and ‘no-dependencies’ are implied.
  • The macros ‘AM_MAINTAINER_MODE’ and ‘AM_CYGWIN32’ are required.

• The check target doesn’t depend on all.

GNU maintainers are advised to use ‘gnu’ strictness in preference to the special Cygnus mode. Some day, perhaps, the differences between Cygnus trees and GNU trees will disappear (for instance, as GCC is made more standards compliant). At that time the special Cygnus mode will be removed.

23 When Automake Isn’t Enough

Automake’s implicit copying semantics means that many problems can be worked around by simply adding some make targets and rules to Makefile.in. Automake will ignore these additions.

There are some caveats to doing this. Although you can overload a target already used by Automake, it is often inadvisable, particularly in the topmost directory of a package with subdirectories. However, various useful targets have a ‘-local’ version you can specify in your Makefile.in. Automake will supplement the standard target with these user-supplied targets.

The targets that support a local version are all, info, dvi, ps, pdf, check,

install-data, install-exec, uninstall, installdirs, installcheck and the various
clean targets (mostlyclean, clean, distclean, and maintainer-clean). Note that
there are no uninstall-exec-local or uninstall-data-local targets; just use

uninstall-local. It doesn’t make sense to uninstall just For instance, here is one way to install a file in /etc:

     install-data-local:
             $(INSTALL_DATA) $(srcdir)/afile $(DESTDIR)/etc/afile

Some targets also have a way to run another target, called a hook, after their work is done. The hook is named after the principal target, with ‘-hook’ appended. The targets

allowing hooks are install-data, install-exec, uninstall, dist, and distcheck.

For instance, here is how to create a hard link to an installed program:

     install-exec-hook:
             ln $(DESTDIR)$(bindir)/program $(DESTDIR)$(bindir)/proglink

24 Distributing Makefile.ins

Automake places no restrictions on the distribution of the resulting Makefile.ins. We still encourage software authors to distribute their work under terms like those of the GPL, but doing so is not required to use Automake.

Some of the files that can be automatically installed via the --add-missing switch do fall under the GPL. However, these also have a special exception allowing you to distribute them with your package, regardless of the licensing you choose.

25 Automake API versioning

New Automake releases usually include bug fixes and new features. Unfortunately they may also introduce new bugs and incompatibilities. This makes four reasons why a package may require a particular Automake version.

Things get worse when maintaining a large tree of packages, each one requiring a different version of Automake. In the past, this meant that any developer (and sometime users) had to install several versions of Automake in different places, and switch ‘$PATH’ appropriately for each package.

Starting with version 1.6, Automake installs versioned binaries. This means you can install several versions of Automake in the same ‘$prefix’, and can select an arbitrary Automake version by running ‘automake-1.6’ or ‘automake-1.7’ without juggling with ‘$PATH’. Furthermore, Makefile’s generated by Automake 1.6 will use ‘automake-1.6’ explicitely in their rebuild rules.

Note that ‘1.6’ in ‘automake-1.6’ is Automake’s API version, not Automake’s version. If a bug fix release is made, for instance Automake 1.6.1, the API version will remain 1.6. This means that a package which work with Automake 1.6 should also work with 1.6.1; after all, this is what people expect from bug fix releases.

Note that if your package relies on a feature or a bug fix introduced in a release, you can pass this version as an option to Automake to ensure older releases will not be used. For instance, use this in your configure.in:

       AM_INIT_AUTOMAKE(1.6.1)    dnl Require Automake 1.6.1 or better.

or, in a particular Makefile.am:

       AUTOMAKE_OPTIONS = 1.6.1   # Require Automake 1.6.1 or better.

Automake will print an error message if its version is older than the requested version.

What is in the API

Automake’s programing interface is not easy to define. Basically it should include at least all documented variables and targets that a ‘Makefile.am’ authors can use, the behaviours associated to them (e.g. the places where ‘-hook’’s are run), the command line interface of

‘automake’ and ‘aclocal’, . . .

What is not in the API

Every undocumented variable, target, or command line option, is not part of the API. You should avoid using them, as they could change from one version to the other (even in bug fix releases, if this helps to fix a bug).

If it turns out you need to use such a undocumented feature, contact automake@gnu. org and try to get it documented and exercised by the test-suite.

Macro and Variable Index 61

Macro and Variable Index

_LDADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 _LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 _LIBADD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 _SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 _TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

A

AC_CANONICAL_HOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CANONICAL_SYSTEM. . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_CONFIG_AUX_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 AC_CONFIG_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 AC_F77_LIBRARY_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . 13 AC_LIBOBJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_LIBSOURCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_LIBSOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_OUTPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 AC_PROG_CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_PROG_F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_PROG_LEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AC_PROG_LIBTOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_PROG_RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_PROG_YACC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AC_SUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 ACLOCAL_AMFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

am_cv_sys_posix_termios . . . . . . . . . . . . . . . . . . . . . 19

AM_C_PROTOTYPES . . . . . . . . . . . . . . . . . . . . . . . 14, 17, 38 AM_CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 AM_CONDITIONAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 AM_CONFIG_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 AM_CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 AM_CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 AM_ETAGSFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 AM_FFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 AM_GCJFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 AM_GNU_GETTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

AM_HEADER_TIOCGWINSZ_NEEDS_SYS_IOCTL . . . . . . 17

AM_INIT_AUTOMAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

AM_INSTALLCHECK_STD_OPTIONS_EXEMPT . . . . . . . . 53

AM_JAVACFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 AM_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 31 AM_MAINTAINER_MODE . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 AM_MAKEINFOFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 AM_PATH_LISPDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 AM_PROG_GCJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 AM_RFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 AM_RUNTESTFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 AUTOCONF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 AUTOMAKE_OPTIONS . . . . . . . . . . . . . . . . . . . . . . 38, 39, 52

B

bin_PROGRAMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 bin_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 build_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 BUILT_SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 CCAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CCASFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 check_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 check_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 check_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 CLASSPATH_ENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 CLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CXXCOMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 CXXLINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

D

data_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 41 DEFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 DEJATOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 DESTDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 dist_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 DIST_SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 49

                                                DISTCHECK_CONFIGURE_FLAGS. . . . . . . . . . . . . . . . . . . 49
                                                distcleancheck_listfiles . . . . . . . . . . . . . . . . 49, 50

DISTCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

                                                distuninstallcheck_listfiles . . . . . . . . . . . . . . . 49

E

ELCFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 ETAGS_ARGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 ETAGSFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 EXPECT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 EXTRA_DIST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 EXTRA_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

F

F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 F77COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 FFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 FLINK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

G

GCJFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 GTAGS_ARGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

H

HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 41 host_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 host_triplet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

I

include_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 INCLUDES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 info_TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

J

JAVA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 JAVAC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 JAVACFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 JAVAROOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

L

LDADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 lib_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 lib_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 LIBADD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 libexec_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 libexec_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 LIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 lisp_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 42 localstate_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

M

MAINTAINERCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . 48 MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MAKEINFO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 MAKEINFOFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 man_MANS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 MANS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 45 MOSTLYCLEANFILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

N

nodist_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 noinst_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 noinst_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 noinst_LISP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 noinst_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 noinst_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 noinst_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

O

oldinclude_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

P

PACKAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 PACKAGE, directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

PACKAGE, prevent definition . . . . . . . . . . . . . . . . . 18

pkgdata_DATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 pkgdata_SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 pkgdatadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkginclude_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 pkgincludedir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkglib_LIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 pkglib_LTLIBRARIES . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 pkglib_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 pkglibdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkgpyexecdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 pkgpythondir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 4 pyexecdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 PYTHON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 43 PYTHON_EXEC_PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 PYTHON_PLATFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 PYTHON_PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 PYTHON_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 pythondir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

R

RFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 RUNTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 RUNTESTDEFAULTFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . 51 RUNTESTFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

S

sbin_PROGRAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 sbin_SCRIPTS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 SCRIPTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 40 sharedstate_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 SOURCES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 SUFFIXES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 sysconf_DATA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

T

TAGS_DEPENDENCIES . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 target_alias. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 TESTS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 TESTS_ENVIRONMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 TEXINFO_TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 TEXINFOS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 44, 45

V

VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

VERSION, prevent definition . . . . . . . . . . . . . . . . . 18

W

WARNINGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 WITH_DMALLOC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 WITH_REGEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

X

XFAIL_TESTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

Y

YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

General Index 64

General Index

#

## (special Automake comment) . . . . . . . . . . . . . . . 2

–

--acdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --add-missing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --cygnus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –enable-debug, example . . . . . . . . . . . . . . . . . . . . . . . . 56

--enable-maintainer-mode . . . . . . . . . . . . . . . . . . . . 14

--force-missing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –gnits, complete description . . . . . . . . . . . . . . . . . . . . 57 --gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 –gnu, complete description . . . . . . . . . . . . . . . . . . . . . 57 –gnu, required files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 15 --include-deps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --no-force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --output-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --print-ac-dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --verbose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 15 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 15 --warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --with-dmalloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --with-regex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 -a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -hook targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -local targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

DATA primary, defined . . . . . . . . . . . . . . . . . . . . . . . 41 DEPENDENCIES, defined . . . . . . . . . . . . . . . . . . . . 24 HEADERS primary, defined. . . . . . . . . . . . . . . . . . . 41 JAVA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 42 LDFLAGS, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 LIBADD primary, defined . . . . . . . . . . . . . . . . . . . . . 26 LIBRARIES primary, defined . . . . . . . . . . . . . . . . . 26 LISP primary, defined. . . . . . . . . . . . . . . . . . . . . . . . . 42 LTLIBRARIES primary, defined . . . . . . . . . . . . . . 27 MANS primary, defined . . . . . . . . . . . . . . . . . . . . . . . 45 PROGRAMS primary variable . . . . . . . . . . . . . . . . . 3 PYTHON primary, defined. . . . . . . . . . . . . . . . . . . . 43 SCRIPTS primary, defined . . . . . . . . . . . . . . . . . . . . 40 SOURCES and header files. . . . . . . . . . . . . . . . . . . . 24 SOURCES primary, defined . . . . . . . . . . . . . . . . . . . 24 TEXINFOS primary, defined . . . . . . . . . . . . . . . . . . 44

@

@ALLOCA@, special handling . . . . . . . . . . . . . . . . . 30 @LIBOBJS@, special handling . . . . . . . . . . . . . . . . . 30 @LTLIBOBJS@, special handling . . . . . . . . . . . . . . 27

A

AC OUTPUT, scanning . . . . . . . . . . . . . . . . . . . . . . . 12 AC_SUBST and SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . 22 acinclude.m4, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 aclocal program, introduction . . . . . . . . . . . . . . . . . . . 6 aclocal search path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 aclocal, extending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 aclocal, Invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 aclocal, Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 aclocal.m4, preexisting . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Adding new SUFFIXES. . . . . . . . . . . . . . . . . . . . . . . . 55 all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 all-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 AM_CONDITIONAL and SUBDIRS . . . . . . . . . . . . . . . . . . 21 AM INIT AUTOMAKE, example use . . . . . . . . . . . 6 ansi2knr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Append operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Automake constraints . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Automake options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Automake requirements . . . . . . . . . . . . . . . . . . . . . 1, 12 Automake, invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Automake, recursive operation . . . . . . . . . . . . . . . . . . 2 Automatic dependency tracking . . . . . . . . . . . . . . . . 39 Automatic linker selection. . . . . . . . . . . . . . . . . . . . . . 37 Auxiliary programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Avoiding path stripping . . . . . . . . . . . . . . . . . . . . . . . . 22

B

BUGS, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 BUILT SOURCES, defined . . . . . . . . . . . . . . . . . . . . 41

C

C++ support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 canonicalizing Automake variables. . . . . . . . . . . . . . . 4 cfortran. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 check primary prefix, definition. . . . . . . . . . . . . . . . . . 4 check-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 check LTLIBRARIES, not allowed . . . . . . . . . . . . . 27 clean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 clean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Comment, special to Automake . . . . . . . . . . . . . . . . . 2 Complete example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Conditional example, –enable-debug . . . . . . . . . . . . 56 Conditional subdirectories. . . . . . . . . . . . . . . . . . . . . . 21 Conditional SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 config.guess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 configure.in, from GNU Hello . . . . . . . . . . . . . . . . . . . 6 configure.in, scanning . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Constraints of Automake . . . . . . . . . . . . . . . . . . . . . . . . 1 cpio example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 cvs-dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 cvs-dist, non-standard example . . . . . . . . . . . . . . . . . . 1 Cygnus strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

D

DATA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 41 de-ANSI-fication, defined. . . . . . . . . . . . . . . . . . . . . . . 38 dejagnu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 depcomp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Dependency tracking. . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Dependency tracking, disabling. . . . . . . . . . . . . . . . . 39 dirlist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Disabling dependency tracking . . . . . . . . . . . . . . . . . 39 dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 dist-bzip2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dist-gzip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 dist-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 59 dist-shar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dist-tarZ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dist-zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dist and nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 DIST_SUBDIRS, explained . . . . . . . . . . . . . . . . . . . . . . . 22 distcheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 distclean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 distclean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 distcleancheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 dmalloc, support for . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 dvi-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

E

E-mail, bug reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 EDITION Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . . 44 else . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 endif. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Example conditional –enable-debug . . . . . . . . . . . . 56 Example of recursive operation . . . . . . . . . . . . . . . . . . 2 Example of shared libraries . . . . . . . . . . . . . . . . . . . . 27 Example, EXTRA PROGRAMS . . . . . . . . . . . . . . . . 3 Example, false and true . . . . . . . . . . . . . . . . . . . . . . . . . 8 Example, GNU Hello. . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Example, handling Texinfo files . . . . . . . . . . . . . . . . . 7 Example, mixed language . . . . . . . . . . . . . . . . . . . . . . 36 Example, regression test. . . . . . . . . . . . . . . . . . . . . . . . . 7 Executable extension. . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Exit status 77, special interpretation . . . . . . . . . . . 51 Expected test failure . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Extending aclocal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Extending list of installation directories . . . . . . . . . 3 Extension, executable . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Extra files distributed with Automake . . . . . . . . . . . 9 EXTRA , prepending . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 EXTRA prog SOURCES, defined . . . . . . . . . . . . . . 25 EXTRA PROGRAMS, defined . . . . . . . . . . . . . . 3, 26

F

false Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Files distributed with Automake . . . . . . . . . . . . . . . . 9 First line of Makefile.am . . . . . . . . . . . . . . . . . . . . . . . . 2 FLIBS, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 foreign strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Fortran 77 support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 Fortran 77, mixing with C and C++ . . . . . . . . . . . . 35 Fortran 77, Preprocessing . . . . . . . . . . . . . . . . . . . . . . 34

G

Gettext support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 gnits strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 gnu strictness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 GNU Gettext support. . . . . . . . . . . . . . . . . . . . . . . . . . 42 GNU Hello, configure.in . . . . . . . . . . . . . . . . . . . . . . . . . 6 GNU Hello, example . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 GNU make extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . 1 GNU Makefile standards . . . . . . . . . . . . . . . . . . . . . . . . 1

H

Header files in SOURCES . . . . . . . . . . . . . . . . . . . . . 24 HEADERS primary, defined. . . . . . . . . . . . . . . . . . . . 41 HEADERS, installation directories . . . . . . . . . . . . . 41 Hello example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Hello, configure.in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 hook targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 HP-UX 10, lex problems . . . . . . . . . . . . . . . . . . . . . . . 19 HTML support, example . . . . . . . . . . . . . . . . . . . . . . . . 3

I

id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 INCLUDES, example usage . . . . . . . . . . . . . . . . . . . . . 7 Including Makefile fragment . . . . . . . . . . . . . . . . . . . . 55 info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 58 info-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 58 Install hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Install, two parts of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 install-data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 install-data-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 install-data-local . . . . . . . . . . . . . . . . . . . . . . . 47, 58 install-exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 58 install-exec-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 install-exec-local . . . . . . . . . . . . . . . . . . . . . . . 47, 58 install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 53 install-info target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 install-man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 53 install-man target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 install-strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Installation directories, extending list . . . . . . . . . . . . 3 Installation support . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 installcheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 installcheck-local . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 installdirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 58 installdirs-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Installing headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Installing scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Invoking aclocal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Invoking Automake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

J

Java support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 JAVA primary, defined . . . . . . . . . . . . . . . . . . . . . . . . . 42 JAVA restrictions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

L

lex problems with HP-UX 10 . . . . . . . . . . . . . . . . . . . 19 lex, multiple lexers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 LIBADD primary, defined . . . . . . . . . . . . . . . . . . . . . . 26 LIBRARIES primary, defined . . . . . . . . . . . . . . . . . . 26 Linking Fortran 77 with C and C++ . . . . . . . . . . . . 35 LISP primary, defined. . . . . . . . . . . . . . . . . . . . . . . . . . 42 local targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 LTLIBRARIES primary, defined . . . . . . . . . . . . . . . 27

M

Macro search path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Macros Automake recognizes . . . . . . . . . . . . . . . . . . . 12 make check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 make clean support . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 make dist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 make distcheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 make distcleancheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 make distuninstallcheck . . . . . . . . . . . . . . . . . . . . . . . . 49 make install support . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 make installcheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Make targets, overriding . . . . . . . . . . . . . . . . . . . . . . . . 2 Makefile fragment, including . . . . . . . . . . . . . . . . . . . 55 Makefile.am, first line . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 MANS primary, defined . . . . . . . . . . . . . . . . . . . . . . . . 45 mdate-sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Mixed language example . . . . . . . . . . . . . . . . . . . . . . . 36 Mixing Fortran 77 with C and C++ . . . . . . . . . . . . . 35 Mixing Fortran 77 with C and/or C++. . . . . . . . . . 35 mostlyclean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 mostlyclean-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Multiple configure.in files. . . . . . . . . . . . . . . . . . . . . . . . 9 Multiple lex lexers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Multiple yacc parsers . . . . . . . . . . . . . . . . . . . . . . . . . . 32

N

no-dependencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 no-installinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 no-installman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 no-texinfo.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 nobase and dist or nodist . . . . . . . . . . . . . . . . . . . 23 nodist and nobase . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 noinst primary prefix, definition . . . . . . . . . . . . . . . . . 4 noinstall-info target . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 noinstall-man target . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Non-GNU packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Non-standard targets. . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

O

Objects in subdirectory . . . . . . . . . . . . . . . . . . . . . . . . 28 Option, ansi2knr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, check-news . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, cygnus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, dejagnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, dist-bzip2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, dist-shar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, dist-tarZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, dist-zip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, no-define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Option, no-dependencies . . . . . . . . . . . . . . . . . . . . . . . 52 Option, no-exeext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, no-installinfo . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, no-installman . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, no-texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, nostdinc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, readme-alpha . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Option, warnings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Options, aclocal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Options, Automake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Options, std-options . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Overriding make targets. . . . . . . . . . . . . . . . . . . . . . . . . 2 Overriding make variables . . . . . . . . . . . . . . . . . . . . . . . 2

P

Path stripping, avoiding. . . . . . . . . . . . . . . . . . . . . . . . 22 pdf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 pdf-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 pkgdatadir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkgincludedir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 pkglibdir, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 POSIX termios headers . . . . . . . . . . . . . . . . . . . . . . . . 19 Preprocessing Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . 34 Primary variable, DATA . . . . . . . . . . . . . . . . . . . . . . . 41 Primary variable, defined. . . . . . . . . . . . . . . . . . . . . . . . 3 Primary variable, HEADERS . . . . . . . . . . . . . . . . . . 41 Primary variable, JAVA . . . . . . . . . . . . . . . . . . . . . . . . 42 Primary variable, LIBADD. . . . . . . . . . . . . . . . . . . . . 26 Primary variable, LIBRARIES . . . . . . . . . . . . . . . . . 26 Primary variable, LISP . . . . . . . . . . . . . . . . . . . . . . . . 42 Primary variable, LTLIBRARIES . . . . . . . . . . . . . . 27 Primary variable, MANS . . . . . . . . . . . . . . . . . . . . . . . 45 Primary variable, PROGRAMS . . . . . . . . . . . . . . . . . 3 Primary variable, PYTHON. . . . . . . . . . . . . . . . . . . . 43 Primary variable, SCRIPTS . . . . . . . . . . . . . . . . . . . . 40 Primary variable, SOURCES . . . . . . . . . . . . . . . . . . . 24 Primary variable, TEXINFOS . . . . . . . . . . . . . . . . . . 44 prog LDADD, defined . . . . . . . . . . . . . . . . . . . . . . . . . 24 Programs, auxiliary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 PROGRAMS primary variable . . . . . . . . . . . . . . . . . . 3 PROGRAMS, bindir . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 ps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 ps-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 PYTHON primary, defined. . . . . . . . . . . . . . . . . . . . . 43

R

Ratfor programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 README-alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Recognized macros by Automake . . . . . . . . . . . . . . . 12 Recursive operation of Automake. . . . . . . . . . . . . . . . 2 regex package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Regression test example . . . . . . . . . . . . . . . . . . . . . . . . . 7 Reporting BUGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Requirements of Automake. . . . . . . . . . . . . . . . . . . . . 12 Requirements, Automake. . . . . . . . . . . . . . . . . . . . . . . . 1 Restrictions for JAVA . . . . . . . . . . . . . . . . . . . . . . . . . . 42 rx package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

S

Scanning configure.in . . . . . . . . . . . . . . . . . . . . . . . . . . 12 SCRIPTS primary, defined . . . . . . . . . . . . . . . . . . . . . 40 SCRIPTS, installation directories . . . . . . . . . . . . . . 41 Selecting the linker automatically . . . . . . . . . . . . . . 37 Shared libraries, support for. . . . . . . . . . . . . . . . . . . . 27 site.exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 SOURCES primary, defined . . . . . . . . . . . . . . . . . . . . 24 Special Automake comment . . . . . . . . . . . . . . . . . . . . . 2 Strictness, command line . . . . . . . . . . . . . . . . . . . . . . . . 9 Strictness, defined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, foreign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, gnits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Strictness, gnu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Subdirectories, building conditionally. . . . . . . . . . . 21 Subdirectory, objects in . . . . . . . . . . . . . . . . . . . . . . . . 28 SUBDIRS and AC_SUBST . . . . . . . . . . . . . . . . . . . . . . . . . 22 SUBDIRS and AM_CONDITIONAL . . . . . . . . . . . . . . . . . . 21 SUBDIRS, conditional . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 SUBDIRS, explained . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 suffix .la, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 suffix .lo, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 SUFFIXES, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Support for C++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Support for Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . . . 34 Support for GNU Gettext . . . . . . . . . . . . . . . . . . . . . . 42 Support for Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

T

tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 TAGS support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Target, install-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Target, install-man. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Target, noinstall-info . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Target, noinstall-man . . . . . . . . . . . . . . . . . . . . . . . . . . 46 termios POSIX headers . . . . . . . . . . . . . . . . . . . . . . . . 19 Test suites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Tests, expected failure . . . . . . . . . . . . . . . . . . . . . . . . . 51 Texinfo file handling example. . . . . . . . . . . . . . . . . . . . 7 Texinfo flag, EDITION. . . . . . . . . . . . . . . . . . . . . . . . . 44 Texinfo flag, UPDATED . . . . . . . . . . . . . . . . . . . . . . . 44 Texinfo flag, UPDATED-MONTH. . . . . . . . . . . . . . 44 Texinfo flag, VERSION . . . . . . . . . . . . . . . . . . . . . . . . 44 texinfo.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 TEXINFOS primary, defined . . . . . . . . . . . . . . . . . . . 44 true Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

U

Uniform naming scheme. . . . . . . . . . . . . . . . . . . . . . . . . 3 uninstall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 58 uninstall-hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 uninstall-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 UPDATED Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . 44 UPDATED-MONTH Texinfo flag . . . . . . . . . . . . . . 44 user variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

V

Variables, overriding . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 variables, reserved for the user . . . . . . . . . . . . . . . . . . 4 VERSION Texinfo flag . . . . . . . . . . . . . . . . . . . . . . . . . 44

W

Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

Y

yacc, multiple parsers . . . . . . . . . . . . . . . . . . . . . . . . . . 32 ylwrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

Z

zardoz example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 General ideas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2.1 General Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2.2 Strictness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 The Uniform Naming Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 How derived variables are named . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Variables reserved for the user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.6 Programs automake might require . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Some example packages . . . . . . . . . . . . . . . . . . . . . . . . . 6

3.1 A simple example, start to finish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 A classic program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.3 Building true and false. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Creating a Makefile.in . . . . . . . . . . . . . . . . . . . . . . . . . . 9

5 Scanning configure.in . . . . . . . . . . . . . . . . . . . . . . . . . 12

5.1 Configuration requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.2 Other things Automake recognizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.3 Auto-generating aclocal.m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.4 aclocal options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.5 Macro search path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.5.1 Modifying the macro search path: --acdir. . . . . . . . . . . . . . . . 15 5.5.2 Modifying the macro search path: -I dir . . . . . . . . . . . . . . . . . 16 5.5.3 Modifying the macro search path: dirlist. . . . . . . . . . . . . . . . 16

5.6 Autoconf macros supplied with Automake . . . . . . . . . . . . . . . . . . . . . 17 5.6.1 Public macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.6.2 Private macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

5.7 Writing your own aclocal macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

6 The top-level Makefile.am. . . . . . . . . . . . . . . . . . . . . . 20

6.1 Recursing subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.2 Conditional subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

6.2.1 Conditional subdirectories with AM_CONDITIONAL . . . . . . . . . . 21 6.2.2 Conditional subdirectories with AC_SUBST . . . . . . . . . . . . . . . . . 22 6.2.3 How DIST_SUBDIRS is used . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

7 An Alternative Approach to Subdirectories . . 22 8 Rebuilding Makefiles. . . . . . . . . . . . . . . . . . . . . . . . . . . 23

9 Building Programs and Libraries . . . . . . . . . . . . . 23

9.1 Building a program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 9.1.1 Defining program sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 9.1.2 Linking the program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 9.1.3 Conditional compilation of sources . . . . . . . . . . . . . . . . . . . . . . . . 25

9.1.3.1 Conditional compilation using _LDADD substitutions . . 25 9.1.3.2 Conditional compilation using Automake conditionals . . 25

9.1.4 Conditional compilation of programs . . . . . . . . . . . . . . . . . . . . . . 26 9.2 Building a library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 9.3 Building a Shared Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.4 Program and Library Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.5 Special handling for LIBOBJS and ALLOCA . . . . . . . . . . . . . . . . . . 30 9.6 Variables used when building a program . . . . . . . . . . . . . . . . . . . . . . . 30 9.7 Yacc and Lex support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.8 C++ Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 9.9 Assembly Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 9.10 Fortran 77 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

9.10.1 Preprocessing Fortran 77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.10.2 Compiling Fortran 77 Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.10.3 Mixing Fortran 77 With C and C++. . . . . . . . . . . . . . . . . . . . . 35

9.10.3.1 How the Linker is Chosen . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 9.10.4 Fortran 77 and Autoconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

9.11 Java Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 9.12 Support for Other Languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 9.13 Automatic de-ANSI-fication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 9.14 Automatic dependency tracking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 9.15 Support for executable extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

10 Other Derived Objects . . . . . . . . . . . . . . . . . . . . . . . 40

10.1 Executable Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 10.2 Header files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 10.3 Architecture-independent data files . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 10.4 Built sources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

11 Other GNU Tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

11.1 Emacs Lisp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 11.2 Gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 11.3 Libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 11.4 Java . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 11.5 Python. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

12 Building documentation . . . . . . . . . . . . . . . . . . . . . . 44

12.1 Texinfo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 12.2 Man pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

13 What Gets Installed . . . . . . . . . . . . . . . . . . . . . . . . . . 46

13.1 Basics of installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 13.2 The two parts of install. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 13.3 Extending installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 13.4 Staged installs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 13.5 Rules for the user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

14 What Gets Cleaned. . . . . . . . . . . . . . . . . . . . . . . . . . . 48

15 What Goes in a Distribution . . . . . . . . . . . . . . . . 48

15.1 Basics of distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 15.2 Fine-grained distribution control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.3 The dist hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.4 Checking the distribution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 15.5 The types of distributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

16 Support for test suites. . . . . . . . . . . . . . . . . . . . . . . . 50

16.1 Simple Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 16.2 DejaGNU Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 16.3 Install Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

17 Changing Automake’s Behavior . . . . . . . . . . . . . 52

18 Miscellaneous Rules . . . . . . . . . . . . . . . . . . . . . . . . . . 54

18.1 Interfacing to etags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 18.2 Handling new file extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 18.3 Support for Multilibs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

19 Include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

20 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

21 The effect of --gnu and --gnits . . . . . . . . . . . . . . 57

22 The effect of --cygnus . . . . . . . . . . . . . . . . . . . . . . . . 58

23 When Automake Isn’t Enough. . . . . . . . . . . . . . . 58

24 Distributing Makefile.ins . . . . . . . . . . . . . . . . . . . . 59

25 Automake API versioning. . . . . . . . . . . . . . . . . . . . 59 Macro and Variable Index. . . . . . . . . . . . . . . . . . . . . . . . . 61

General Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64